### v1

# Lecture Notes on High Dimensional Linear Regression

Lecture Notes on High Dimensional Linear Regression
Alberto Quaini111 Department of Econometrics, Erasmus University of Rotterdam, P.O. Box 1738, 3000 DR Rotterdam, The Netherlands. 222 Please, let me know if you find typos or mistakes at quaini@ese.eur.nl.
(December 20, 2024 (first version: October 17, 2024))
Introduction
These lecture notes were developed for a Master’s course in advanced machine learning at Erasmus University of Rotterdam. The course is designed for graduate students in mathematics, statistics and econometrics. The content follows a proposition-proof structure, making it suitable for students seeking a formal and rigorous understanding of the statistical theory underlying machine learning methods.
At present, the notes focus on linear regression, with an in-depth exploration of the existence, uniqueness, relations, computation, and non-asymptotic properties of the most prominent estimators in this setting: least squares, ridgeless, ridge, and lasso.
Background
It is assumed that readers have a solid background in calculus, linear algebra, convex analysis, and probability theory. Some definitions and results from these fields, relevant for the course, are provided in the Appendix for reference.
Book-length references
The content of these lecture notes is inspired by a wide range of existing literature, but the presentation of topics follows my own interpretation and logical structure. Although most of the content can be traced back to established sources, certain sections reflect my perspective, and some material is original to this course. For those interested in more comprehensive, book-length discussions of related topics, the following key references are recommended: Hastie et al. (2009), Bühlmann and Van De Geer (2011), Hastie et al. (2015), and Wainwright (2019).
Disclaimer
Please note that despite my efforts, these lecture notes may contain errors. I welcome any feedback, corrections, or suggestions you may have. If you spot any mistakes or have ideas for improvement, feel free to contact me via email at quaini@ese.eur.nl.
Notation
  • All random variables are defined on complete probability space $(\Omega,\mathcal{F},\mathbb{P})$ and take values in a real Euclidean space.
  • For a random variable (vector) [matrix] $x$ ($\bm{x}$) [$\bm{X}$], the notation $x\in\mathbb{R}$ ($\bm{x}\in\mathbb{R}^{n}$) [$\bm{X}\in\mathbb{R}^{n\times p}$] means that $x$ ($\bm{x}$) [$\bm{X}$] takes values in $\mathbb{R}$ ($\mathbb{R}^{n}$) [$\mathbb{R}^{n\times p}$].
[…]
  • Given a random variable $x$, its expectation is denoted $\mathbb{E}[x]$ and its variance $\operatorname{\mathbb{V}ar}[x]$.
  • For a vector $\bm{x}\in\mathbb{R}^{n}$, the $i-$th element is denoted $x_{i}$ for $i=1,\ldots,n$.
  • For a matrix $\bm{A}\in\mathbb{R}^{n\times p}$, the $i,j-$th element is denoted $A_{i,j}$, the $j-$th column is denoted $\bm{A}_{j}$ and the $i-$th row is denoted $A_{(i)}$, for $i=1,\ldots,n$ and $j=1,\ldots,p$.
  • The transpose of a matrix $\bm{A}\in\mathbb{R}^{n\times p}$ is denoted $\bm{A}^{\prime}$.
  • The Moore-Penrose inverse of a matrix $\bm{A}\in\mathbb{R}^{n\times p}$ is denoted $\bm{A}^{+}$.
  • The $l_{p}-$norm $\left\lVert\cdot\right\rVert_{p}$ on $\mathbb{R}^{n}$ is defined for all $\bm{v}\in\mathbb{R}^{n}$ as $\left\lVert\bm{v}\right\rVert_{p}:=\left(\sum_{i=1}^{n}|v_{i}|^{p}\right)^{1/p}$ when $p\in[1,+\infty)$, and $\left\lVert\bm{v}\right\rVert_{p}:=\max_{i=1}^{n}|v_{i}|$ when $p=+\infty$.
  • Given vector $\bm{\theta}\in\mathbb{R}^{p}$, the $l_{0}-$norm (which is not a norm!) $\left\lVert\bm{x}\right\rVert_{0}$ counts the number of nonzero elements of $\bm{\theta}$.
  • $\operatorname*{argmin}_{x\in X}f(x)$ denotes the set of minimizers of $f$ over set $X$.
  • $\operatorname{diag}(\bm{A})$ denotes the diagonal elements of a square matrix $\bm{A}\in\mathbb{R}^{n\times n}$
  • $\operatorname{diag}(a_{1},\ldots,a_{n})$ denotes a square matrix in $\mathbb{R}^{n\times n}$ that has diagonal elements given by $a_{1},\ldots,a_{n}\in\mathbb{R}$ and that has zero elsewhere.
  • Given a matrix $\bm{A}\in\mathbb{R}^{n\times p}$, its rank is $\operatorname{Rank}(\bm{A})$, its range is $\operatorname{Range}(\bm{A})$, its kernel is $\operatorname{Ker}(\bm{A})$, its trace is denoted $\operatorname{Trace}(\bm{A})$.
  • We denote $\operatorname{P}_{S}$ the orthogonal projection onto set $S\subset\mathbb{R}^{n\times p}$.
  • Given a vector space $V$, the sum of two subsets $A,B\subset V$ is defined as $A+B:=\{a+b\ :\ a\in A,b\in B\}$. The sum of a set $A\in V$ and a vector $b\in V$ is defined as $A+b:=\{a+b\ :\ a\in A\}$.
  • The symbol $\partial$ indicates the subdifferential.
Chapter 1 Linear Regression
Linear regression is a supervised learning technique aimed at predicting a target random variable $y$ using a linear combination
$\bm{x}^{\prime}\bm{\theta}=x_{1}\theta_{1}+\ldots+x_{p}\theta_{p}$
of explanatory variables $\bm{x}=[x_{1},\ldots,x_{p}]^{\prime}$, where $\bm{\theta}\in\mathbb{R}^{p}$ and $p\in\mathbb{N}$111 An intercept in $f(\bm{x};\bm{\theta})$ can be introduced by adding a constant term to the predictors. The target variable $y$ is also referred to as dependent or output variable, while the explanatory variables $\bm{x}$ are also known as independent variables, predictors or input variables.
In typical applications, we observe only a sample of size $n\in\mathbb{N}$ of these random variables, represented by the pairs $(\bm{x}_{i},y_{i})_{i=1}^{n}$, where $\bm{x}_{i}\in\mathbb{R}^{p}$ and $y_{i}\in\mathbb{R}$ for each $i$. Given a regression coefficient $\bm{\theta}_{0}\in\mathbb{R}^{p}$, a statistical linear model, or simply linear model, is expressed as
$y_{i}=\bm{x}_{i}^{\prime}\bm{\theta}_{0}+\varepsilon_{0i},\quad i=1,\ldots,n,$ (1.1)
where $\varepsilon_{0i}$ are real-valued residual random variables. Figure 1.1 depicts a linear model for $i=1,\ldots,n$ observations $y_{i}$, with two predictors $\tilde{\bm{x}}_{i}=[1,x_{i}]^{\prime}$ consisting of a unit constant and a variable $x_{i}$, a coefficient $\bm{\theta}_{0}\in\mathbb{R}^{2}$ and the associated error terms $\varepsilon_{0i}$. The Data Generating Process (DGP), i.e., the joint distribution of the predictors $\bm{x}$ and the real-valued residual random variables $\bm{\varepsilon}_{0}=[\varepsilon_{0i}]_{i=1}^{n}$, is subject to certain restrictions. Depending on the type of restrictions imposed on the DGP, different types of linear models are obtained. The two general forms of linear models are fixed and random design models, which are defined as follows.
Definition 1 (Fixed design model).
In a fixed design model, the sequence $(\bm{x}_{i})_{i=1}^{n}$ is fixed. The residuals $\varepsilon_{0i}$ are independent and identically distributed.
Definition 2 (Random design model).
In a random design model, the pair $(\bm{x}_{i},y_{i})_{i=1}^{n}$ is a sequence of independent and identically distributed random variables.
The fixed design model is particularly suitable when the predictors are controlled by the analyst, such as the dose of medication administered to patients in the treatment group in a clinical trial. Conversely, the random design model is appropriate when the explanatory variables are stochastic, such as the wind speed observed at a specific time and location.
Figure 1.1: Statistical linear model $y_{i}=\tilde{\bm{x}}_{i}^{\prime}\bm{\theta}_{0}+\varepsilon_{0i}$ where $\tilde{\bm{x}}_{i}=[1,x_{i}]^{\prime}$ and $\bm{\theta}_{0}=[0.5,0.8]^{\prime}$.
We organize the observed values of the target variable in the vector $\bm{y}=[y_{1},\ldots,y_{n}]^{\prime}\in\mathbb{R}^{n}$, and the observations on the predictors in the design matrix
$\bm{X}:=\begin{bmatrix}x_{11}&\ldots&x_{1p}\\ \vdots&\ddots&\vdots\\ x_{n1}&\ldots&x_{np}\end{bmatrix}\mathbb{R}^{n\times p}.$
With this notation, the linear model (1.1) can be expressed as:
$\bm{y}=\bm{X}\bm{\theta}_{0}+\bm{\varepsilon}_{0},$
where $\bm{\varepsilon}_{0}=[\varepsilon_{01},\ldots,\varepsilon_{0n}]^{\prime}\in\mathbb{R}^{n}$.
Example 1.
A classic example of linear regression is found in the work of Nerlove et al. (1961), which examines returns to scale in the U.S. electricity power supply industry. In this study, the total cost $y_{i}$ for firm $i$ is predicted using a linear model based on the firm’s output production $x_{i1}$, the wage rate $x_{i2}$, the price of fuel $x_{i3}$, and the rental price of capital $x_{i4}$, with data from $n=145$ electric utility companies.
The rest of the chapter is organized as follows. First, we study the most basic linear regression approach, the method of least squares projection, and some of its recent machine learning extensions. Our study focuses on their existence, uniqueness, connections, geometric interpretation, and computation. Then, we study we cover both their finite- or small-sample properties, that are valid for any given sample size, and their asymptotic properties, that are useful approximations when the sample size is large enough.
1.0.1 Estimators and their properties
Definition 3 (Estimand).
An estimand is a feature, or parameter, of interest of the population.
Definition 4 (estimator and estimate).
An estimator is a function taking as input the data, and possibly other auxiliary variables, and outputting an estimate, which is a specific value assigned to the estimand.
For instance, in the context of the linear model (1.1), the coefficient $\bm{\theta}_{0}\in\mathbb{R}^{p}$ represents the estimand. An estimator is a function $\hat{\bm{\theta}}_{n}:\mathbb{R}^{n}\times\mathbb{R}^{n\times p}\to\mathbb{R}^{p}$ that takes as inputs the data $(\bm{y},\bm{X})\in\mathbb{R}^{n}\times\mathbb{R}^{n\times p}$ and produces an estimate $\hat{\bm{\theta}}_{n}(\bm{y},\bm{X})\in\mathbb{R}^{p}$. For simplicity, we use the same notation, $\hat{\bm{\theta}}_{n}$, to refer to both the estimator and the resulting estimate, although formally the estimate should be written as $\hat{\bm{\theta}}_{n}(\bm{y},\bm{X})$. We may also write $\hat{\bm{\theta}}_{n}\in\mathbb{R}^{p}$ to indicate that an estimator outputs values in $\mathbb{R}^{p}$.
Definition 5 (Linear prediction).
The quantity $\operatorname{lm}(\bm{X},\bm{\theta}):=\bm{X}\bm{\theta}$ denotes the linear prediction associated to coefficient vector $\bm{\theta}\in\mathbb{R}^{p}$ and predictors $\bm{X}\in\mathbb{R}^{n\times p}$.
Let $\bm{M}^{+}$ denote the Moore-Penrose inverse of a generic real-valued matrix $\bm{M}$. We make extensive use of the following useful projections.222 We use the notation $\operatorname{Range}$ ($\operatorname{Ker}$) for the range (kernel) of a matrix. Details on these sets, the Moore-Penrose inverse and orthogonal projections are given in Appendix Section LABEL:app:sec:linear_algebra.
Definition 6 (Useful projections).
Given a fixed matrix $\bm{X}\in\mathbb{R}^{n\times p}$:
  • $\bm{\operatorname{P}}_{\operatorname{Range}(\bm{X}^{\prime})}:=\bm{X}^{+}\bm{X}$ is the orthogonal projector onto $\operatorname{Range}(\bm{X}^{\prime})$;
  • $\bm{\operatorname{P}}_{\operatorname{Ker}(\bm{X})}:=\bm{I}-\bm{X}^{+}\bm{X}$ is the orthogonal projector onto $\operatorname{Ker}(\bm{X})$;
  • $\bm{\operatorname{P}}_{\operatorname{Range}(\bm{X})}:=\bm{X}\bm{X}^{+}$ is the orthogonal projector onto $\operatorname{Range}(\bm{X})$;
  • $\bm{\operatorname{P}}_{\operatorname{Ker}(\bm{X}^{\prime})}:=\bm{I}-\bm{X}\bm{X}^{+}$ is the orthogonal projector onto $\operatorname{Ker}(\bm{X}^{\prime})$.
The next proposition demonstrates that, if we fix the design matrix $\bm{X}$, we can focus on regression coefficients in $\operatorname{Range}(\bm{X}^{\prime})$. Indeed, coefficients in this set span all possible linear predictions that can be achieved through $\bm{X}$.
Proposition 1.0.1.
Given a matrix $\bm{X}\in\mathbb{R}^{n\times p}$, for any $\bm{\theta}\in\mathbb{R}^{p}$:
$\operatorname{lm}(\bm{X},\bm{\theta})=\operatorname{lm}(\bm{X},\bm{\operatorname{P}}_{\operatorname{Range}(\bm{X}^{\prime})}\bm{\theta}).$
Proof.
Using the identity $\bm{I}=\bm{\operatorname{P}}_{\operatorname{Range}(\bm{X}^{\prime})}+\bm{\operatorname{P}}_{\operatorname{Ker}(\bm{X})}$, we have for any $\bm{\theta}\in\mathbb{R}^{p}$:
$\bm{X}\bm{\theta}=\bm{X}(\bm{\operatorname{P}}_{\operatorname{Range}(\bm{X}^{\prime})}+\bm{\operatorname{P}}_{\operatorname{Ker}(\bm{X})})\bm{\theta}=\bm{X}\bm{\operatorname{P}}_{\operatorname{Range}(\bm{X}^{\prime})}\bm{\theta}.$
Finite-sample properties
Since an estimator is derived from data, it is a random variable. Intuitively, when comparing two estimators of the same estimand, we prefer the one whose probability distribution is ”more concentrated” around the true value of the estimand. Formally, estimators are compared using several key properties.333 Some of these properties are defined by means of the $l_{2}-$norm. Note that this choice is typical, but arbitrary.
Definition 7 (Bias).
The bias of an estimator $\hat{\bm{\theta}}_{n}$ of $\bm{\theta}_{0}$ is the difference between the expected value of the estimator and the estimand:
$\operatorname{Bias}(\hat{\bm{\theta}}_{n},\bm{\theta}_{0}):=\mathbb{E}[\hat{\bm{\theta}}_{n}]-\bm{\theta}_{0}.$
Definition 8 (Estimation risk).
The estimation risk of an estimator $\hat{\bm{\theta}}_{n}$ of $\bm{\theta}_{0}$ measures the difference between the estimate and the estimand as:
$\operatorname{ER}(\hat{\bm{\theta}}_{n},\bm{\theta}_{0}):=\lVert\hat{\bm{\theta}}_{n}-\bm{\theta}_{0}\rVert_{2}^{2}.$
Definition 9 (MSE).
The Mean Squared Error (MSE) of an estimator $\hat{\bm{\theta}}_{n}$ of $\bm{\theta}_{0}$ is the expected estimation risk of the estimator:
$\operatorname{MSE}(\hat{\bm{\theta}}_{n},\bm{\theta}_{0}):=\mathbb{E}[\lVert\hat{\bm{\theta}}_{n}-\bm{\theta}_{0}\rVert_{2}^{2}].$
Definition 10 (Predictive risk).
The predictive risk of an estimator $\hat{\bm{\theta}}_{n}$ of $\bm{\theta}_{0}$ measures the difference between the linear predictions of $\hat{\bm{\theta}}_{n}$ and those of $\bm{\theta}_{0}$:
$\operatorname{PR}(\hat{\bm{\theta}}_{n},\bm{\theta}_{0}):=\lVert\operatorname{lm}(\bm{X},\hat{\bm{\theta}}_{n})-\operatorname{lm}(\bm{X},\bm{\theta}_{0})\rVert_{2}^{2}/n.$
Definition 11 (Mean predictive risk).
The Mean Predictive Risk (MPR) of an estimator $\hat{\bm{\theta}}_{n}$ of $\bm{\theta}_{0}$ is the expected predictive risk of the estimator:
$\operatorname{MPR}(\hat{\bm{\theta}}_{n},\bm{\theta}_{0}):=\mathbb{E}[\lVert\operatorname{lm}(\bm{X},\hat{\bm{\theta}}_{n})-\operatorname{lm}(\bm{X},\bm{\theta}_{0})\rVert_{2}^{2}/n].$
As a corollary to Proposition 1.0.1, the predictive risk of an estimator is unchanged if both the estimator and the estimand are projected onto $\operatorname{Range}(\bm{X}^{\prime})$.
Corollary 1.0.1.
Given a matrix $\bm{X}\in\mathbb{R}^{n\times p}$, for any estimator $\hat{\bm{\theta}}_{n}$ of $\bm{\theta}_{0}\in\mathbb{R}^{p}$:
$\operatorname{PR}(\hat{\bm{\theta}}_{n},\bm{\theta}_{0})=\operatorname{PR}(\bm{\operatorname{P}}_{\operatorname{Range}(\bm{X}^{\prime})}\hat{\bm{\theta}}_{n},\bm{\operatorname{P}}_{\operatorname{Range}(\bm{X}^{\prime})}\bm{\theta}_{0}).$
Proof.
The result follows from Proposition 1.0.1. ∎
The next proposition justifies the definition of mean predictive risk given in Definition 11.
Proposition 1.0.2.
Assume that the linear model (1.1) holds with $\mathbb{E}[\bm{x}\varepsilon_{0}]=\bm{0}$. Then, for any $\bm{\theta}\in\mathbb{R}^{p}$:
$\mathbb{E}[\lVert\bm{y}-\bm{X}\bm{\theta}\rVert_{2}^{2}/n]=\operatorname{MPR}(\bm{\theta},\bm{\theta}_{0})+\mathbb{E}[\left\lVert\bm{\varepsilon}_{0}\right\rVert_{2}^{2}/n].$
Proof.
Since $\bm{y}=\bm{X}\bm{\theta}_{0}+\bm{\varepsilon}_{0}$, we have
$\mathbb{E}[\lVert\bm{y}-\bm{X}\bm{\theta}\rVert_{2}^{2}/n]=\mathbb{E}[\lVert\bm{X}(\bm{\theta}_{0}-\bm{\theta})\rVert_{2}^{2}/n]+\mathbb{E}[\left\lVert\bm{\varepsilon}_{0}\right\rVert_{2}^{2}/n]+2(\bm{\theta}_{0}-\bm{\theta})\mathbb{E}[\bm{X}^{\prime}\bm{\varepsilon}_{0}].$
Then, the result follows since $\mathbb{E}[\bm{X}^{\prime}\bm{\varepsilon}_{0}]=\sum_{i=1}^{n}\mathbb{E}[\bm{X}_{i}\varepsilon_{0i}]=\bm{0}$, where $\bm{X}_{i}$ denotes the $i-$th row of $\bm{X}$. ∎
If our primary goal is to accurately predict the target variable, we seek a estimators $\hat{\bm{\theta}}$ with a low mean prediction risk $\mathbb{E}[\lVert\bm{y}-\bm{X}\hat{\bm{\theta}}\rVert_{2}^{2}/n]$. Since we cannot control the error term $\varepsilon_{0}$, Proposition 1.0.2 suggests that we should focus on estimators with a low mean predictive risk.
On the other hand, if our interest lies in understanding which predictors influence the target variable and how they do so, the true coefficient $\bm{\theta}_{0}$ becomes our focus. In this case, we might prefer unbiased estimators – those with zero bias – over biased ones. However, estimators with lower mean squared error (MSE) are generally favored, even if they feature some bias. The following proposition demonstrates that the MSE can be decomposed into a bias and a variance term.
Proposition 1.0.3 (Bias-variance decomposition of MSE).
Given an estimator $\hat{\bm{\theta}}_{n}\in\mathbb{R}^{p}$ for $\bm{\theta}_{0}\in\mathbb{R}^{p}$, the MSE can be decomposed as follows:
$\operatorname{MSE}(\hat{\bm{\theta}}_{n},\bm{\theta}_{0})=\lVert\operatorname{Bias}(\hat{\bm{\theta}}_{n},\bm{\theta}_{0})\rVert_{2}^{2}+\operatorname{Trace}(\operatorname{\mathbb{V}ar}[\hat{\bm{\theta}}_{n}]).$
Proof.
The result follows from
$\operatorname{MSE}(\hat{\bm{\theta}}_{n},\bm{\theta}_{0})$ $=\mathbb{E}[(\hat{\bm{\theta}}_{n}-\bm{\theta}_{0})^{\prime}(\hat{\bm{\theta}}_{n}-\bm{\theta}_{0})]$
$=\mathbb{E}[\operatorname{Trace}\{(\hat{\bm{\theta}}_{n}-\bm{\theta}_{0})(\hat{\bm{\theta}}_{n}-\bm{\theta}_{0})^{\prime}\}]$
$=\operatorname{Trace}(\mathbb{E}[(\hat{\bm{\theta}}_{n}-\bm{\theta}_{0})(\hat{\bm{\theta}}_{n}-\bm{\theta}_{0})^{\prime}])$
$=\operatorname{Trace}(\mathbb{E}[(\hat{\bm{\theta}}_{n}-\mathbb{E}[\hat{\bm{\theta}}_{n}]+\operatorname{Bias}(\hat{\bm{\theta}}_{n},\bm{\theta}_{0}))(\hat{\bm{\theta}}_{n}-\mathbb{E}[\hat{\bm{\theta}}_{n}]+\operatorname{Bias}(\hat{\bm{\theta}}_{n},\bm{\theta}_{0}))^{\prime}])$
$=\operatorname{Trace}(\operatorname{\mathbb{V}ar}[\hat{\bm{\theta}}_{n}]+\operatorname{Bias}(\hat{\bm{\theta}}_{n},\bm{\theta}_{0})\operatorname{Bias}(\hat{\bm{\theta}}_{n},\bm{\theta}_{0})^{\prime}+$
$\qquad\qquad\ \mathbb{E}[\hat{\bm{\theta}}_{n}-\mathbb{E}[\hat{\bm{\theta}}_{n}]]\operatorname{Bias}(\hat{\bm{\theta}}_{n},\bm{\theta}_{0})^{\prime}+\operatorname{Bias}(\hat{\bm{\theta}}_{n},\bm{\theta}_{0})\mathbb{E}[\hat{\bm{\theta}}_{n}-\mathbb{E}[\hat{\bm{\theta}}_{n}]]^{\prime})$
$=\operatorname{Trace}(\operatorname{\mathbb{V}ar}[\hat{\bm{\theta}}_{n}]+\operatorname{Bias}(\hat{\bm{\theta}}_{n},\bm{\theta}_{0})\operatorname{Bias}(\hat{\bm{\theta}}_{n},\bm{\theta}_{0})^{\prime})$
$=\lVert\operatorname{Bias}(\hat{\bm{\theta}}_{n},\bm{\theta}_{0})\rVert_{2}^{2}+\operatorname{Trace}(\operatorname{\mathbb{V}ar}[\hat{\bm{\theta}}_{n}]).$
Loosely speaking, the bias and the variance of an estimator are linked to the estimator’s ”complexity”. Estimators with higher complexity often fit the data better, resulting in lower bias, but they are more sensitive to data variations, leading to higher variance. Conversely, estimators with lower complexity tend to have lower variance but higher bias, a phenomenon known as the bias-variance tradeoff.
Apart from simple cases, computing the finite-sample properties of estimators, such as their MSE or predictive risk, is infeasible or overly complicated. This is because they require computations under the DGP of complex transformations of the data. When direct computation is not possible, we can rely on concentration inequalities or asymptotic approximations.
Concentration inequalities are inequalities that bound the probability that a random variable deviates from a particular value, typically its expectation. In this chapter, we focus on inequalities that control the MSE or predictive risk of an estimator, such as:
$\mathbb{P}[d(\hat{\bm{\theta}}_{n},\bm{\theta}_{0})\leq h(\bm{y},\bm{X},n,p)]\geq 1-\delta,$
or
$\mathbb{P}[d(\operatorname{lm}(\bm{X},\hat{\bm{\theta}}_{n}),\operatorname{lm}(\bm{X},\bm{\theta}_{0}))\leq h(\bm{y},\bm{X},n,p)]\geq 1-\delta,$
where $\delta\in(0,1)$ is the level of confidence, $d:\mathbb{R}^{p}\times\mathbb{R}^{p}\to[0,+\infty)$ is a distance, and $h$ is a real-valued function of the data, the sample size, and the number of predictors.
Large-sample properties
Large-sample or asymptotic theory provides an alternative approach to study and analyse estimators. Classically, this framework develops approximations of the finite-sample properties of estimators, such as their distribution, MSE or predictive risk, by letting the sample size $n\to\infty$. Consequently, these approximations work well when the sample size $n$ is much larger than the number of predictors $p$. More recently, asymptotic approximations are also developed by letting $p\to\infty$, or having both $n,p\to\infty$ at some rate. Note that, given a sample of size $n$ and number of variables $p$, there is no general indication on how to choose the appropriate asymptotic regime for $n$ and $p$, as the goodness of fit of the corresponding asymptotic approximations should be assessed on a case by case basis. In this chapter, we work with two notions from large-sample theory: consistency and asymptotic distribution.
Definition 12 (Consistency).
Estimator $\hat{\bm{\theta}}_{n}$ of $\bm{\theta}_{0}$ is consistent, written $\hat{\bm{\theta}}_{n}\overset{\mathbb{P}}{\to}\bm{\theta}_{0}$ as $n\to\infty$, if for all $\varepsilon>0$,
$\lim_{n\to\infty}\mathbb{P}[\lvert\hat{\bm{\theta}}_{n}-\bm{\theta}_{0}\rvert>\varepsilon]=0.$
Definition 13 (Asymptotic distribution).
[…]

### v2

@@
 Lecture Notes on High Dimensional Linear Regression
-Alberto Quaini111 Department of Econometrics, Erasmus University of Rotterdam, P.O.
+Alberto Quaini [^1] Department of Econometrics, Erasmus University of Rotterdam, P.O.
 Box 1738, 3000 DR Rotterdam, The Netherlands.
-222 Please, let me know if you find typos or mistakes at quaini@ese.eur.nl.
-(December 20, 2024 (first version: October 17, 2024))
+[^2] Please, let me know if you find typos or mistakes at quaini@ese.eur.nl.
+(December 3, 2025 (first version: October 17, 2024))
 Introduction
@@
 The content follows a proposition-proof structure, making it suitable for students seeking a formal and rigorous understanding of the statistical theory underlying machine learning methods.
-At present, the notes focus on linear regression, with an in-depth exploration of the existence, uniqueness, relations, computation, and non-asymptotic properties of the most prominent estimators in this setting: least squares, ridgeless, ridge, and lasso.
+At present, the notes focus on linear regression, with an in-depth exploration of the existence, uniqueness, relations, computation, and non-asymptotic properties of prominent estimators in this setting: least squares, ridgeless, ridge, and lasso.
 Background
-It is assumed that readers have a solid background in calculus, linear algebra, convex analysis, and probability theory.
+It is assumed that readers have a solid background in mathematical analysis, linear algebra, convex analysis, and probability theory.
 Some definitions and results from these fields, relevant for the course, are provided in the Appendix for reference.
@@
   • All random variables are defined on complete probability space $(\Omega,\mathcal{F},\mathbb{P})$ and take values in a real Euclidean space.
-  • For a random variable (vector) [matrix] $x$ ($\bm{x}$) [$\bm{X}$], the notation $x\in\mathbb{R}$ ($\bm{x}\in\mathbb{R}^{n}$) [$\bm{X}\in\mathbb{R}^{n\times p}$] means that $x$ ($\bm{x}$) [$\bm{X}$] takes values in $\mathbb{R}$ ($\mathbb{R}^{n}$) [$\mathbb{R}^{n\times p}$].
+  • For a random variable (vector) [matrix] $x$ ($\bm{x}$) $\bm{X}$], the notation $x\in\mathbb{R}$ ($\bm{x}\in\mathbb{R}^{n}$) $\bm{X}\in\mathbb{R}^{n\times p}$] means that $x$ ($\bm{x}$) $\bm{X}$] takes values in $\mathbb{R}$ ($\mathbb{R}^{n}$) $\mathbb{R}^{n\times p}$].
 […]
@@
 $\bm{x}^{\prime}\bm{\theta}=x_{1}\theta_{1}+\ldots+x_{p}\theta_{p}$
-of explanatory variables $\bm{x}=[x_{1},\ldots,x_{p}]^{\prime}$, where $\bm{\theta}\in\mathbb{R}^{p}$ and $p\in\mathbb{N}$111 An intercept in $f(\bm{x};\bm{\theta})$ can be introduced by adding a constant term to the predictors.
+of explanatory variables $\bm{x}=[x_{1},\ldots,x_{p}]^{\prime}$, where $\bm{\theta}\in\mathbb{R}^{p}$ and $p\in\mathbb{N}$.
+[^1] An intercept in $f(\bm{x};\bm{\theta})$ can be introduced by adding a constant term to the predictors.
 The target variable $y$ is also referred to as dependent or output variable, while the explanatory variables $\bm{x}$ are also known as independent variables, predictors or input variables.
@@
 We organize the observed values of the target variable in the vector $\bm{y}=[y_{1},\ldots,y_{n}]^{\prime}\in\mathbb{R}^{n}$, and the observations on the predictors in the design matrix
-$\bm{X}:=\begin{bmatrix}x_{11}&\ldots&x_{1p}\\ \vdots&\ddots&\vdots\\ x_{n1}&\ldots&x_{np}\end{bmatrix}\mathbb{R}^{n\times p}.$
+$\bm{X}:=\begin{bmatrix}x_{11}&\ldots&x_{1p}\\ \vdots&\ddots&\vdots\\ x_{n1}&\ldots&x_{np}\end{bmatrix}\in\mathbb{R}^{n\times p}.$
 With this notation, the linear model (1.1) can be expressed as:
@@
 Let $\bm{M}^{+}$ denote the Moore-Penrose inverse of a generic real-valued matrix $\bm{M}$.
-We make extensive use of the following useful projections.222 We use the notation $\operatorname{Range}$ ($\operatorname{Ker}$) for the range (kernel) of a matrix.
+We make extensive use of the following useful projections.
+[^2] We use the notation $\operatorname{Range}$ ($\operatorname{Ker}$) for the range (kernel) of a matrix.
 Details on these sets, the Moore-Penrose inverse and orthogonal projections are given in Appendix Section LABEL:app:sec:linear_algebra.
@@
   • $\bm{\operatorname{P}}_{\operatorname{Ker}(\bm{X}^{\prime})}:=\bm{I}-\bm{X}\bm{X}^{+}$ is the orthogonal projector onto $\operatorname{Ker}(\bm{X}^{\prime})$.
-The next proposition demonstrates that, if we fix the design matrix $\bm{X}$, we can focus on regression coefficients in $\operatorname{Range}(\bm{X}^{\prime})$.
-Indeed, coefficients in this set span all possible linear predictions that can be achieved through $\bm{X}$.
+The next proposition demonstrates that, if we fix the design matrix $\bm{X}$, we can focus on regression coefficients in $\operatorname{Range}(\bm{X}^{\prime})$. Indeed,
+$\mathbb{R}^{p}=\operatorname{Ker}(\bm{X})+\operatorname{Range}(\bm{X}^{\prime})$
+and coefficients in $\operatorname{Ker}(\bm{X})$ are immaterial for linear prediction, while coefficients in $\operatorname{Range}(\bm{X}^{\prime})$ span all possible linear predictions that can be achieved through $\bm{X}$.
 Proposition 1.0.1.
@@
 Intuitively, when comparing two estimators of the same estimand, we prefer the one whose probability distribution is ”more concentrated” around the true value of the estimand.
-Formally, estimators are compared using several key properties.333 Some of these properties are defined by means of the $l_{2}-$norm.
+Formally, estimators are compared using several key properties.
+[^3] Some of these properties are defined by means of the $l_{2}$-norm.
 Note that this choice is typical, but arbitrary.
@@
 Definition 8 (Estimation risk).
-The estimation risk of an estimator $\hat{\bm{\theta}}_{n}$ of $\bm{\theta}_{0}$ measures the difference between the estimate and the estimand as:
+The estimation risk of an estimator $\hat{\bm{\theta}}_{n}$ of $\bm{\theta}_{0}$ measures the difference between the estimator and the estimand as:
 $\operatorname{ER}(\hat{\bm{\theta}}_{n},\bm{\theta}_{0}):=\lVert\hat{\bm{\theta}}_{n}-\bm{\theta}_{0}\rVert_{2}^{2}.$
@@
 $\operatorname{MPR}(\hat{\bm{\theta}}_{n},\bm{\theta}_{0}):=\mathbb{E}[\lVert\operatorname{lm}(\bm{X},\hat{\bm{\theta}}_{n})-\operatorname{lm}(\bm{X},\bm{\theta}_{0})\rVert_{2}^{2}/n].$
+While the MSE and and the MPR are deterministic quantities, the estimation risk and the predictive risk are stochastic, unless we treat the design matrix as fixed.
 As a corollary to Proposition 1.0.1, the predictive risk of an estimator is unchanged if both the estimator and the estimand are projected onto $\operatorname{Range}(\bm{X}^{\prime})$.
@@
 Proof.
-The result follows from
+The result follows from $\mathbb{E}[\hat{\bm{\theta}}_{n}-\mathbb{E}[\hat{\bm{\theta}}_{n}]]=0$ and
 $\operatorname{MSE}(\hat{\bm{\theta}}_{n},\bm{\theta}_{0})$ $=\mathbb{E}[(\hat{\bm{\theta}}_{n}-\bm{\theta}_{0})^{\prime}(\hat{\bm{\theta}}_{n}-\bm{\theta}_{0})]$
@@
 Definition 13 (Asymptotic distribution).
+Given a deterministic real-valued sequence $r_{n,p}\to\infty$, let $F_{n,p}$ be the probability distribution of $r_{n,p}(\hat{\bm{\theta}}_{n}-\bm{\theta}_{0})$ and $F$ a non-degenerate probability distribution.
+Estimator $\hat{\bm{\theta}}_{n}$ of $\bm{\theta}_{0}$ has asymptotic distribution $F$ with rate of convergence $r_{n,p}$ if $F_{n,p}(\bm{z})\to F(\bm{z})$ as $r_{n,p}\to\infty$ for all $\bm{z}$ at which $F(\bm{z})$ is continuous.
+Equivalent short-hand notations are $r_{n,p}(\hat{\bm{\theta}}_{n}-\bm{\theta}_{0})\overset{d}{\to}F$ and $r_{n,p}(\hat{\bm{\theta}}_{n}-\bm{\theta}_{0})\overset{d}{\to}\bm{\eta}\sim F$, as $r_{n,p}\to\infty$.
+1.1 Least Squares and Penalized Least Squares
+In this chapter, we study the most widely used methods in linear regression analysis: the method of least squares and some of its penalized variants.
+The method of least squares was first introduced by Legendre (1805) and Gauss (1809), and it consists in minimizing the squared $l_{2}$-distance between the target values $\bm{y}$ and the linear prediction $\operatorname{lm}(\bm{X},\bm{\theta})=\bm{X}\bm{\theta}$ in the coefficient vector $\bm{\theta}\in\mathbb{R}^{p}$.
+Figure 1.2: Adrien-Marie Legendre (1752–1833) and Johann Carl Friedrich Gauss (1977–1855).
+Definition 14 (Least squares estimator).
+The Least Squares Estimator (LSE) is defined as:
+$\hat{\bm{\theta}}_{n}^{ls}\in\operatorname*{argmin}_{\bm{\theta}\in\mathbb{R}^{p}}\frac{1}{2}\sum_{i=1}^{n}(y_{i}-\bm{\theta}^{\prime}\bm{x}_{i})^{2}=\operatorname*{argmin}_{\bm{\theta}\in\mathbb{R}^{p}}\frac{1}{2}\left\lVert\bm{y}-\bm{X}\bm{\theta}\right\rVert_{2}^{2}.$ (1.2)
+In addition to the LSE, we consider the following variants.
+Definition 15 (Ridgeless estimator).
+The ridgeless estimator is defined as:
+$\hat{\bm{\theta}}_{n}^{rl}=\operatorname*{argmin}_{\hat{\bm{\theta}}\in\mathbb{R}^{p}}\left\{\lVert\hat{\bm{\theta}}\rVert_{2}^{2}\ :\ \hat{\bm{\theta}}\in\operatorname*{argmin}_{\bm{\theta}\in\mathbb{R}^{p}}\frac{1}{2}\left\lVert\bm{y}-\bm{X}\bm{\theta}\right\rVert_{2}^{2}\right\}.$ (1.3)
+Definition 16 (Ridge estimator).
+The ridge estimator is defined for $\lambda>0$ as:
+$\hat{\bm{\theta}}_{n}^{r}(\lambda)=\operatorname*{argmin}_{\bm{\theta}\in\mathbb{R}^{p}}\frac{1}{2}\left\lVert\bm{y}-\bm{X}\bm{\theta}\right\rVert_{2}^{2}+\frac{\lambda}{2}\left\lVert\bm{\theta}\right\rVert_{2}^{2}.$ (1.4)
+Definition 17 (Lasso estimator).
+The lasso estimator is defined for $\lambda>0$ as:
+$\hat{\bm{\theta}}_{n}^{l}(\lambda)\in\operatorname*{argmin}_{\bm{\theta}\in\mathbb{R}^{p}}\frac{1}{2}\left\lVert\bm{y}-\bm{X}\bm{\theta}\right\rVert_{2}^{2}+\lambda\left\lVert\bm{\theta}\right\rVert_{1}.$ (1.5)
+Here is a brief overview of the results that are discussed in detail in the rest of this chapter.
+A solution to the least squares problem (1.2) always exists.
+However, when the predictors (i.e., the columns of $\bm{X}$) are linearly dependent, there are infinitely many solutions.
+[^4] This situation always arises when $p>n$, and it may arise even when $n\leq p$.
+In such cases, the LSE typically considered is the ridgeless estimator, which is always unique.
+The ridge and lasso estimators are penalized or regularized versions of the LSE, with penalty term $\lambda\left\lVert\bm{\theta}\right\rVert_{2}^{2}$ and $\lambda\left\lVert\bm{\theta}\right\rVert_{1}$, respectively.
+The penalty parameter $\lambda>0$ controls the strength of the penalty.
+The ridge estimator, introduced by Hoerl and Kennard (1970), was developed to address certain shortcomings of the LSE, particularly in scenarios involving collinear or multicollinear designs – where the predictors in $\bm{X}$ are linearly dependent or nearly-linearly dependent.
+The ridge estimator is uniquely defined and often exhibits better statistical properties compared to those of the LSE in settings with multicollinear or many predictors.
+On the other hand, the lasso estimator, popularized by Tibshirani (1996), offers an approximation of the $l_{0}$ estimator, which is defined for some $R>0$:
+$\hat{\bm{\theta}}_{n}^{l_{0}}(\lambda)\in\operatorname*{argmin}_{\bm{\theta}\in\mathbb{R}^{p}}\left\{\frac{1}{2}\left\lVert\bm{y}-\bm{X}\bm{\theta}\right\rVert_{2}^{2}\ :\ \left\lVert\bm{\theta}\right\rVert_{0}\leq R\right\},$ (1.6)
+where $\left\lVert\bm{\theta}\right\rVert_{0}$ is the number of nonzero elements in $\bm{\theta}$.
+A key feature of this estimator is its ability to produce sparse solutions, i.e., to set some coefficients exactly to zero.
+Consequently, the $l_{0}$ estimator can be used to perform parameter estimation and variable selection simultaneously.
+However, it is the solution of a non-convex problem, and, in general, computing it can be an ”NP-hard” problem.
+The lasso instead shares the ability to produce sparse solutions and it can be easily computed even for large datasets.
+Remark 1 (Data standardization).
+For computational stability, it is recommended to compute linear regression estimators with a least squares loss after having standardized the predictors $\bm{X}$ so that $\bar{\bm{x}}:=\bm{X}^{\prime}\bm{1}/n=\bm{0}$ and $\bm{X}_{j}^{\prime}\bm{X}_{j}=1$ for each $j=1\ldots,p$.
+Without standardization, the solutions would depend on the units used to measure the predictors.
+Moreover, we may also center the target variable $\bm{y}$, meaning $\bar{y}:=\bm{y}^{\prime}\bm{1}/n=0$.
+These centering conditions are convenient, since they mean that we can omit the intercept term.
+Given an optimal solution $\hat{\bm{\theta}}$ on the centered data, we can recover the optimal solutions for the uncentered data: $\hat{\bm{\theta}}$ is the same and the intercept is given by $\bar{y}-\bar{\bm{x}}^{\prime}\hat{\bm{\theta}}$.
+1.1.1 Existence and uniqueness
+From here on, we make extensive use of the spectral decomposition of $\bm{X}$.
+Definition 18 (Spectral decomposition of $\bm{X}$).
+The spectral decomposition of $\bm{X}$ is
+$\bm{X}=\bm{U}\bm{S}\bm{V}^{\prime},$
+where $\bm{U}\in\mathbb{R}^{n\times n}$ and $\bm{V}\in\mathbb{R}^{p\times p}$ are orthogonal matrices and, for $r:=\operatorname{Rank}(\bm{X})\leq\min\{n,p\},$
+$\bm{S}=\begin{bmatrix}\operatorname{diag}(s_{1},\ldots,s_{r})&\bm{0}\\ \bm{0}&\bm{0}\end{bmatrix}\in\mathbb{R}^{n\times p}.$
+We establish the following key results: the existence of the LSE, the ridgeless, the ridge and the lasso estimators; the closed-form expression of the LSE, ridgeless, and ridge; the uniqueness of the ridgeless and ridge; the uniqueness of the LSE when $\operatorname{Rank}(\bm{X})=p$, i.e., when the predictors in $\bm{X}$ are linearly independent.
+Notice that this rank condition cannot hold if $n<p$.
+Theorem 1.1.1 (Existence and uniqueness of LSE, ridgeless, ridge and lasso).
+  1. The following statements hold:
+  2. (i)
+    The set of solutions to the least squares problem (1.2) is non-empty and given by
+    $\operatorname*{argmin}_{\bm{\theta}\in\mathbb{R}^{p}}\left\lVert\bm{y}-\bm{X}\bm{\theta}\right\rVert_{2}^{2}/2=\bm{X}^{+}\bm{y}+\operatorname{Ker}(\bm{X}).$
+  3. (ii)
+    The ridgeless estimator exists, is an element of $\operatorname{Range}(\bm{X}^{\prime})$, and is uniquely given in closed form by
+    $\hat{\bm{\theta}}_{n}^{rl}=\bm{X}^{+}\bm{y}.$ (1.7)
+  4. (iii)
+    If $\operatorname{Rank}(\bm{X})=p$, then the LSE and the ridgeless estimator are uniquely given in closed form by:
+    $\hat{\bm{\theta}}_{n}^{ls}=\hat{\bm{\theta}}_{n}^{rl}=(\bm{X}^{\prime}\bm{X})^{-1}\bm{X}^{\prime}\bm{y}.$ (1.8)
+  5. (iv)
+    The ridge estimator with penalty parameter $\lambda>0$ exists, is an element of $\operatorname{Range}(\bm{X}^{\prime})$, and is uniquely given in closed form by:
+    $\hat{\bm{\theta}}_{n}^{r}(\lambda)=(\bm{X}^{\prime}\bm{X}+\lambda\bm{I})^{-1}\bm{X}^{\prime}\bm{y}.$ (1.9)
+  6. (v)
+    The lasso estimator exists and, in general, it is not unique.
+Proof.
+  1. (i)
+    The least squares problem (1.2) is an unconstrained optimization problem with the objective function
+    $f:\mathbb{R}^{p}\to[0,+\infty);\ \bm{\theta}\mapsto\left\lVert\bm{y}-\bm{X}\bm{\theta}\right\rVert_{2}^{2}/2.$
+    By Theorem LABEL:thm:_fermat's_rule, the set of least squares minimizers is given by
+    $S:=\{\hat{\bm{\theta}}\in\mathbb{R}^{p}:\bm{X}^{\prime}\bm{X}\hat{\bm{\theta}}=\bm{X}^{\prime}\bm{y}\}.$
+    Since $\bm{X}^{\prime}\bm{y}\in\operatorname{Range}(\bm{X}^{\prime})$ and $\operatorname{Range}(\bm{X}^{\prime})=\operatorname{Range}(\bm{X}^{\prime}\bm{X})$, set $S$ is not empty.
+Consider a vector $\hat{\bm{\theta}}\in\bm{X}^{+}\bm{y}+\operatorname{Ker}(\bm{X})$.
+Using that $\bm{X}=\bm{\operatorname{P}}_{\operatorname{Range}(\bm{X})}\bm{X}$ which implies $\bm{X}^{\prime}=\bm{X}^{\prime}\bm{\operatorname{P}}_{\operatorname{Range}(\bm{X})}$, we obtain
+    $\bm{X}^{\prime}\bm{X}\hat{\bm{\theta}}=\bm{X}^{\prime}\bm{X}\bm{X}^{+}\bm{y}=\bm{X}^{\prime}\bm{\operatorname{P}}_{\operatorname{Range}(\bm{X})}\bm{y}=\bm{X}^{\prime}\bm{y}.$
+    Therefore, $\bm{X}^{+}\bm{y}+\operatorname{Ker}(\bm{X})\subset S$.
+Now consider a vector $\bm{v}\in\mathbb{R}^{p}$ not in set $\bm{X}^{+}\bm{y}+\operatorname{Ker}(\bm{X})$.
+That is, $\bm{v}=\hat{\bm{\theta}}+\bm{u}$ with $\hat{\bm{\theta}}\in\bm{X}^{+}\bm{y}+\operatorname{Ker}(\bm{X})$ and $\bm{u}\in\operatorname{Range}(\bm{X}^{\prime})$.
+Since $\bm{X}\bm{u}\not=\bm{0}$,
+    $\bm{X}^{\prime}\bm{X}\bm{v}=\bm{X}^{\prime}\bm{y}+\bm{X}^{\prime}\bm{X}\bm{u}\not=\bm{X}^{\prime}\bm{y}.$
+    We conclude that $\bm{X}^{+}\bm{y}+\operatorname{Ker}(\bm{X})=S$.
+  2. (ii)
+    The minimum norm least squares problem in (1.3) has a strictly convex and coercive objective function
+    $f:\mathbb{R}^{p}\to\mathbb{R};\bm{\theta}\mapsto\left\lVert\bm{\theta}\right\rVert_{2}^{2},$
+    and a closed convex feasible set $\bm{X}^{+}\bm{y}+\operatorname{Ker}(\bm{X})\subset\mathbb{R}^{p}$.
+It follows that a solution exists and it is unique; see Propositions LABEL:app:prop:existence_minimizers and LABEL:app:prop:uniqueness_minimizers.
+Since $\bm{X}^{+}\bm{y}\in\operatorname{Range}(\bm{X}^{\prime})$ and $\bm{v}\in\operatorname{Ker}(\bm{X})=\operatorname{Range}(\bm{X}^{\prime})^{\perp}$ are orthogonal, we have
+    $\|\bm{X}^{+}\bm{y}\|_{2}^{2}\leq\|\bm{X}^{+}\bm{y}\|_{2}^{2}+\|\bm{v}\|_{2}^{2}\leq\|\bm{X}^{+}\bm{y}+\bm{v}\|_{2}^{2},$
+    and therefore
+    $\|\bm{X}^{+}\bm{y}\|_{2}\leq\|\bm{X}^{+}\bm{y}+\bm{v}\|_{2}.$
+    Thus, the ridgeless estimator can be expressed in closed form as $\hat{\bm{\theta}}_{n}^{rl}=\bm{X}^{+}\bm{y}$, which is an element of $\operatorname{Range}(\bm{X}^{\prime})$ since $\bm{X}^{+}=\bm{\operatorname{P}}_{\operatorname{Range}(\bm{X}^{\prime})}\bm{X}^{+}$.
+  3. (iii)
+    If $\operatorname{Rank}(\bm{X})=p$, then $\operatorname{Ker}(\bm{X})=\{\bm{0}\}$.
+Moreover, $\bm{X}^{\prime}\bm{X}$ is invertible and we can use the identity $\bm{X}^{+}=(\bm{X}^{\prime}\bm{X})^{-1}\bm{X}^{\prime}$ to conclude that the LSE and the ridgeless estimator are uniquely given by (1.8).
+  4. (iv)
+    The ridge problem in (1.4) is an unconstrained optimization problem with the strictly convex, coercive and continuously differentiable objective function
+    $f:\mathbb{R}^{p}\to[0,+\infty);\ \bm{\theta}\mapsto\left\lVert\bm{y}-\bm{X}^{\prime}\bm{\theta}\right\rVert_{2}^{2}/2+\lambda/2\left\lVert\bm{\theta}\right\rVert_{2}^{2}.$
+    It follows that a solution $\hat{\bm{\theta}}^{r}\in\mathbb{R}^{n}$ exists and it is unique; see, again, Propositions LABEL:app:prop:existence_minimizers and LABEL:app:prop:uniqueness_minimizers.
+Theorem LABEL:thm:_fermat's_rule implies
+    $(\bm{X}^{\prime}\bm{X}+\lambda\bm{I})\hat{\bm{\theta}}^{r}(\lambda)=\bm{X}^{\prime}\bm{y}.$
+    Consider the spectral decomposition $\bm{X}=\bm{U}\bm{S}\bm{V}^{\prime}$ in Definition 18. Then,
+    $\bm{X}^{\prime}\bm{X}+\lambda\bm{I}=\bm{V}\bm{S}^{\prime}\bm{S}\bm{V}^{\prime}+\lambda\bm{V}\bm{V}^{\prime}=\bm{V}\begin{bmatrix}\operatorname{diag}(s_{1}^{2},\ldots,s_{r}^{2})+\lambda&\bm{0}\\ \bm{0}&\lambda\bm{I}\end{bmatrix}\bm{V}^{\prime},$
+    which is positive definite, and thus $\hat{\bm{\theta}}_{n}^{r}(\lambda)=(\bm{X}^{\prime}\bm{X}+\lambda\bm{I})^{-1}\bm{X}^{\prime}\bm{y}$ is the solution to the FOCs.
+Finally, to prove that $\hat{\bm{\theta}}_{n}^{r}(\lambda)\in\operatorname{Range}(\bm{X}^{\prime})$, notice that $\bm{\operatorname{P}}_{\operatorname{Range}(\bm{X}^{\prime})}=\bm{V}\bm{S}^{+}\bm{S}\bm{V}^{\prime}$, where
+    $\bm{S}^{+}\bm{S}=\begin{bmatrix}\bm{I}&\bm{0}\\ \bm{0}&\bm{0}\end{bmatrix}$.
+    Thus,
+    $\bm{\operatorname{P}}_{\operatorname{Range}(\bm{X}^{\prime})}\hat{\bm{\theta}}_{n}^{r}(\lambda)=$ $\bm{V}\bm{S}^{+}\bm{S}\bm{V}^{\prime}\bm{V}(\bm{S}^{\prime}\bm{S}+\lambda\bm{I})^{-1}\bm{V}^{\prime}\bm{V}\bm{S}^{\prime}\bm{U}^{\prime}\bm{y}$
+    $=$ $\bm{V}(\bm{S}^{\prime}\bm{S}+\lambda\bm{I})^{-1}\bm{S}^{\prime}\bm{U}^{\prime}\bm{y}=\hat{\bm{\theta}}_{n}^{r}(\lambda).$
+    We conclude that $\bm{\operatorname{P}}_{\operatorname{Range}(\bm{X}^{\prime})}\hat{\bm{\theta}}_{n}^{r}(\lambda)=\hat{\bm{\theta}}_{n}^{r}(\lambda)$, i.e., $\hat{\bm{\theta}}_{n}^{r}(\lambda)\in\operatorname{Range}(\bm{X}^{\prime})$.
+  5. (v)
+    The lasso problem in (1.5) is an unconstrained optimization problem with convex and coercive objective function
+    $f:\mathbb{R}^{p}\to[0,+\infty),\qquad\bm{\theta}\mapsto\tfrac{1}{2}\|\bm{y}-\bm{X}\bm{\theta}\|_{2}^{2}+\lambda\|\bm{\theta}\|_{1}.$
+    Hence a solution $\hat{\bm{\theta}}^{l}(\lambda)\in\mathbb{R}^{p}$ always exists; see Proposition LABEL:app:prop:existence_minimizers.
+To see that such a solution is not necessarily unique, consider the case $(\bm{y},\bm{X})\in\mathbb{R}^{n}\times\mathbb{R}^{n\times 2}$ where the two predictors coincide, i.e.
+$\bm{x}_{1}=\bm{x}_{2}=\bm{x}\in\mathbb{R}^{n}$.
+For any $(\theta_{1},\theta_{2})\in\mathbb{R}^{2}$, the lasso objective reduces to
+    $f(\theta_{1},\theta_{2})=\tfrac{1}{2}\|\bm{y}-\bm{x}(\theta_{1}+\theta_{2})\|_{2}^{2}+\lambda(|\theta_{1}|+|\theta_{2}|).$
+    Writing $s=\theta_{1}+\theta_{2}$, we have $|\theta_{1}|+|\theta_{2}|\geq|s|$ with equality if $\theta_{1}$ and $\theta_{2}$ have the same sign.
+Hence the minimization problem reduces to the univariate lasso in $s$:
+    $\min_{s\in\mathbb{R}}\;\tfrac{1}{2}\|\bm{y}-\bm{x}s\|_{2}^{2}+\lambda|s|.$
+    Let $s^{\star}$ denote the optimal solution of this problem.
+Then any coefficient vector of the form $s^{\star},0]^{\prime}$ or $0,s^{\star}]^{\prime}$ (and more generally, any decomposition $\theta_{1},\theta_{2}]^{\prime}$ with $\theta_{1}+\theta_{2}=s^{\star}$ and $|\theta_{1}|+|\theta_{2}|=|s^{\star}|$) achieves the same objective value.
+These are distinct whenever $s^{\star}\neq 0$, establishing that the lasso solution is not unique in this example.
+Remark 2 (Computation ridgeless and ridge).
+The closed form expressions of the LSE, ridgeless and ridge estimators are useful analytical result.
+However, for numerical stability, it is recommended to compute these estimators by solving their corresponding normal equations, which are $\bm{X}^{\prime}\bm{X}\hat{\bm{\theta}}_{n}=\bm{X}^{\prime}\bm{y}$ for the LSE or ridgless, and $(\bm{X}^{\prime}\bm{X}+\lambda\bm{I})\hat{\bm{\theta}}_{n}^{r}=\bm{X}^{\prime}\bm{y}$ for the ridge.
+Remark 3 (Collinearity).
+Using the notation in Definition 18, the minimum nonzero eigenvalue of $\bm{X}^{\prime}\bm{X}$ is $s_{r}^{2}$.
+If $r<p$, then $\bm{X}^{\prime}\bm{X}$ has prp-r zero eigenvalues and the predictors are said to be collinear, that is, they are linearly dependent.
+In this case $\operatorname{Ker}(\bm{X})$ is not trivial (it contains nonzero elements), hence the LSE is not unique.
+Moreover, if $s_{r}\approx 0$, then the computation of
+$\bm{X}^{+}=\bm{V}\begin{bmatrix}\operatorname{diag}(1/s_{1},\ldots,1/s_{r})&\bm{0}\\ \bm{0}&\bm{0}\end{bmatrix}\bm{U}^{\prime},$
+and hence of the ridgeless estimator, is unstable.
+The ridge estimator instead may not display these computational hurdles, provided that the penalty parameter $\lambda$ is large enough.
+That is because the minimum eigenvalue of $(\bm{X}^{\prime}\bm{X}+\lambda I)$ is $s_{r}^{2}+\lambda$.
+In Section 1.1.5 we show that, if $s_{r}\approx 0$, the ridgeless (ridge) estimator’s MSE and MPR satisfy loose (sharp) concentration inequalities.
+Remark 4 (Uniqueness of the lasso solution).
+Tibshirani (2013) shows that, under some conditions, the lasso estimator is unique.
+For instance, if the predictors in $\bm{X}$ are in general position, then the lasso solution is unique.
+Specifically, a set $(\bm{a}_{j})_{j=1}^{p}$ where $\bm{a}_{j}\in\mathbb{R}^{n}$ for all $j$ is in general position if any affine subspace of $\mathbb{R}^{n}$ of dimension $k<n$ contains at most $k+1$ elements of the set $\{\pm\bm{x}_{1},\pm\bm{x}_{2},...\pm\bm{x}_{p}\}$, excluding antipodal pairs of points (that is, points differing only by a sign flip).
+If the predictors are (non redundant) continuous random variables, they are almost surely in general position, and hence the lasso solution is unique.
+As a result, non-uniqueness of the lasso solution typically occurs with discrete-valued data, such as those comprising dummy or categorical variables.
+Since the LSE, ridgeless, ridge and lasso estimators exist, their linear predictions exist too.
+Moreover, the linear predictions of the uniquely defined estimators, like ridgeless and ridge, are trivially unique.
+Remarkably, some estimators that may not be unique entail unique linear predictions.
+The next lemma implies that the LSE and lasso are among these estimators.
+Lemma 1.1.2.
+Let $h:\mathbb{R}^{p}\to(-\infty,+\infty]$ be a proper convex function.
+Then $\bm{X}\bm{\theta}_{1}=\bm{X}\bm{\theta}_{2}$ and $h(\bm{\theta}_{1})=h(\bm{\theta}_{2})$ for every minimizers $\bm{\theta}_{1},\bm{\theta}_{2}\in\mathbb{R}^{p}$ of
+$f:\mathbb{R}^{p}\to(-\infty,+\infty];\bm{\theta}\mapsto\frac{1}{2}\left\lVert\bm{y}-\bm{X}\bm{\theta}\right\rVert_{2}^{2}+h(\bm{\theta}).$
+Proof.
+Assume that $\bm{X}\bm{\theta}_{1}\neq\bm{X}\bm{\theta}_{2}$, and let $\delta:=\inf_{\bm{\theta}\in\mathbb{R}^{p}}f(\bm{\theta})$.
+By Proposition LABEL:app:prop:argmin_convex, the set of minimizers of $f$ is convex.
+Thus, for any $\alpha\in(0,1)$:
+$\delta=$ $f(\alpha\bm{\theta}_{1}+(1-\alpha)\bm{\theta}_{2})$
+$=$ $\frac{1}{2}\left\lVert\bm{y}-\bm{X}[\alpha\bm{\theta}_{1}+(1-\alpha)\bm{\theta}_{2}]\right\rVert_{2}^{2}+h(\alpha\bm{\theta}_{1}+(1-\alpha)\bm{\theta}_{2})$
+$<$ $\alpha\frac{1}{2}\left\lVert\bm{y}-\bm{X}\bm{\theta}_{1}\right\rVert_{2}^{2}+(1-\alpha)\frac{1}{2}\left\lVert\bm{y}-\bm{X}\bm{\theta}_{2}\right\rVert_{2}^{2}+h(\alpha\bm{\theta}_{1}+(1-\alpha)\bm{\theta}_{2})$
+$\leq$ $\alpha\frac{1}{2}\left\lVert\bm{y}-\bm{X}\bm{\theta}_{1}\right\rVert_{2}^{2}+(1-\alpha)\frac{1}{2}\left\lVert\bm{y}-\bm{X}\bm{\theta}_{2}\right\rVert_{2}^{2}+\alpha h(\bm{\theta}_{1})+(1-\alpha)h(\bm{\theta}_{2})$
+$=$ $\alpha f(\bm{\theta}_{1})+(1-\alpha)f(\bm{\theta}_{2})=\delta,$
+where the strict inequality follows from the strict convexity of $g:\bm{v}\mapsto\left\lVert\bm{y}-\bm{v}\right\rVert_{2}^{2}$.
+Since the conclusion $\delta<\delta$ is absurd, we must have $\bm{X}\bm{\theta}_{1}=\bm{X}\bm{\theta}_{2}$; and since $f(\bm{\theta}_{1})=f(\bm{\theta}_{2})$, it follows that $h(\bm{\theta}_{1})=h(\bm{\theta}_{2})$.
+∎
+While we make use of this lemma for proving uniqueness of the predictions of lasso, we can use a more direct approach for the other estimators’ predictions, which directly provides their closed form expressions.
+We further show that the LSE’s prediction has the geometric interpretation of being the unique vector in the range of $\bm{X}$ that is closest to $\bm{y}$ in $l_{2}$ distance, and that the residual vector is orthogonal to the range of $\bm{X}$.
+Theorem 1.1.3 (Uniqueness of linear predictions).
+The following statements hold:
+  1. (i)
+    The linear predictions of the LSE and the ridgeless estimator are uniquely given by:
+    $\operatorname{lm}(\bm{X},\hat{\bm{\theta}}_{n}^{ls})=\operatorname{lm}(\bm{X},\hat{\bm{\theta}}_{n}^{rl})=\bm{\operatorname{P}}_{\operatorname{Range}(\bm{X})}\bm{y},$ (1.10)
+    which is the unique vector $\bm{v}\in\operatorname{Range}(\bm{X})$ such that
+    $\left\lVert\bm{y}-\bm{v}\right\rVert_{2}=\inf\{\left\lVert\bm{y}-\bm{z}\right\rVert_{2}\ :\ \bm{z}\in\operatorname{Range}(\bm{X})\}.$
+    Moreover, the residual vector $\bm{y}-\operatorname{lm}(\bm{X},\hat{\bm{\theta}}_{n}^{ls})=\bm{y}-\operatorname{lm}(\bm{X},\hat{\bm{\theta}}_{n}^{rl})$ is orthogonal to $\operatorname{Range}(\bm{X})$.
+  2. (ii)
+    The linear prediction of the ridge estimator is uniquely given, for $\lambda>0$, by
+    $\operatorname{lm}(\bm{X},\hat{\bm{\theta}}_{n}^{r}(\lambda))=\bm{X}(\bm{X}^{\prime}\bm{X}+\lambda\bm{I})^{-1}\bm{X}^{\prime}\bm{y}.$
+  3. (iii)
+    The linear prediction of the lasso estimator is unique.
+Proof.
+  1. (i)
+    The linear predictions $\operatorname{lm}(\bm{X},\hat{\bm{\theta}}^{ls}_{n})$ and $\operatorname{lm}(\bm{X},\hat{\bm{\theta}}_{n}^{rl})$ are uniquely given by (1.10) because all solutions to the least squares problem $\hat{\bm{\theta}}\in\bm{X}^{+}\bm{y}+\operatorname{Ker}(\bm{X})$ yield the same prediction
+    $\operatorname{lm}(\bm{X},\hat{\bm{\theta}})=\bm{X}\bm{X}^{+}\bm{y}=\bm{\operatorname{P}}_{\operatorname{Range}(\bm{X})}\bm{y}.$
+    By the definition of $\hat{\bm{\theta}}_{n}^{ls}$ and the fact that $\operatorname{Range}(\bm{X})$ is a closed vector subspace of $\mathbb{R}^{n}$, the remaining claims follow as a direct application of the Hilbert projection theorem (Theorem LABEL:app:thm:hilber_projection_theorem).
+  2. (ii)
+    This result follows directly from the closed form expression (1.9) of the ridge estimator.
+  3. (iii)
+    Since the $l_{1}$-norm is convex, the result follows by Lemma 1.1.2.
+1.1.2 Equivalent expressions and relations
+The ridgeless and the ridge, together with their corresponding linear predictions, admit the following simple expressions.
+Proposition 1.1.1 (Spectral expression of ridgeless and ridge).
+Given the spectral decomposition $\bm{X}=\bm{U}\bm{S}\bm{V}^{\prime}$ in Definition 18:
+  1. (i)
+    The ridgeless estimator is given by
+    $\hat{\bm{\theta}}^{rl}_{n}=\left(\sum_{j=1}^{r}\frac{1}{s_{j}}\bm{v}_{j}\bm{u}_{j}^{\prime}\right)\bm{y}.$
+    The corresponding linear prediction is
+    $\operatorname{lm}(\bm{X},\hat{\bm{\theta}}^{rl}_{n})=\left(\sum_{j=1}^{r}\bm{u}_{j}\bm{u}_{j}^{\prime}\right)\bm{y}.$ (1.11)
+  2. (ii)
+    The ridge estimator with $\lambda>0$ is given by
+    $\hat{\bm{\theta}}^{r}_{n}(\lambda)=\left(\sum_{j=1}^{r}\frac{s_{j}}{s_{j}^{2}+\lambda}\bm{v}_{j}\bm{u}_{j}^{\prime}\right)\bm{y}.$
+    The corresponding linear prediction is
+    $\operatorname{lm}(\bm{X},\hat{\bm{\theta}}^{r}_{n})=\left(\sum_{j=1}^{r}\frac{s_{j}^{2}}{s_{j}^{2}+\lambda}\bm{u}_{j}\bm{u}_{j}^{\prime}\right)\bm{y}.$ (1.12)
+Proof.
+  1. (i)
+    From the closed-form expression of the ridgeless estimator,
+    $\hat{\bm{\theta}}^{rl}_{n}=\bm{X}^{+}\bm{y}=\bm{V}\bm{S}^{+}\bm{U}^{\prime}\bm{y}=\left(\sum_{j=1}^{r}\frac{1}{s_{j}}\bm{v}_{j}\bm{u}_{j}^{\prime}\right)\bm{y}.$
+    Therefore,
+    $\bm{X}\hat{\bm{\theta}}^{rl}_{n}=\bm{U}\bm{S}\bm{S}^{+}\bm{U}^{\prime}\bm{y}=\left(\sum_{j=1}^{r}\bm{u}_{j}\bm{u}_{j}^{\prime}\right)\bm{y}.$
+  2. (ii)
+    From the closed-form expression of the ridge estimator,
+    $\hat{\bm{\theta}}^{r}_{n}=(\bm{X}^{\prime}\bm{X}+\lambda\bm{I})^{-1}\bm{X}^{\prime}\bm{y}=$ $\bm{V}(\bm{S}^{\prime}\bm{S}+\lambda\bm{I})^{-1}\bm{S}^{\prime}\bm{U}^{\prime}\bm{y}$
+    $=$ $\left(\sum_{j=1}^{r}\frac{s_{j}}{s_{j}^{2}+\lambda}\bm{v}_{j}\bm{u}_{j}^{\prime}\right)\bm{y}.$
+    Therefore,
+    $\bm{X}\hat{\bm{\theta}}^{r}_{n}=\bm{U}\bm{S}(\bm{S}^{\prime}\bm{S}+\lambda\bm{I})^{-1}\bm{S}^{\prime}\bm{U}^{\prime}\bm{y}=\left(\sum_{j=1}^{r}\frac{s_{j}^{2}}{s_{j}^{2}+\lambda}\bm{u}_{j}\bm{u}_{j}^{\prime}\right)\bm{y}.$
+Using Definition 18, matrix $\bm{\operatorname{P}}_{\operatorname{Range}(\bm{X})}=\bm{X}\bm{X}^{+}=\sum_{j=1}^{r}\bm{u}_{j}\bm{u}_{j}^{\prime}$, where $\{\bm{u}_{1},\ldots,\bm{u}_{r}\}$ is an orthogonal basis of $\operatorname{Range}(\bm{X})$, i.e., they define orthogonal directions in the observation space $\mathbb{R}^{n}$.
+From expression (1.11), it follows that the prediction of the ridgeless estimator is the orthogonal projection of $\bm{y}$ onto the range of $\bm{X}$.
+Expression (1.12) instead shows that the ridge estimator shrinks this projection, shrinking less the directions $\bm{u}_{j}$ associated to high variance (high $s_{j}$), and more the directions $\bm{u}_{j}$ associated to low variance (low $s_{j}$); see Figure 1.3.
+Indeed, for fixed $\lambda>0$, the weight $s_{j}^{2}/(s_{j}^{2}+\lambda)\to 0$ as $s_{j}\to 0$, and $s_{j}^{2}/(s_{j}^{2}+\lambda)\to 1$ as $s_{j}\to\infty$.
+Each direction $\bm{u}_{j}$ corresponds to a latent principal component of the fitted values $\bm{X}\hat{\bm{\theta}}$: the associated singular value $s_{j}$ measures how much variation in the data $\bm{X}$ lies along that direction.
+Hence, directions with large $s_{j}$ correspond to high-variance directions of $\bm{X}$, while directions with small $s_{j}$ correspond to low-variance or nearly collinear directions.
+High-variance directions (large $s_{j}$) are those that $\bm{X}$ can represent accurately; they correspond to well-identified combinations of regressors, and ridge leaves them nearly unshrunk.
+Low-variance directions (small $s_{j}$) are those where $\bm{X}$ has little leverage, so small perturbations in $\bm{y}$ would cause large changes in the least-squares estimate; ridge regularization dampens those directions heavily by multiplying by $s_{j}^{2}/(s_{j}^{2}+\lambda)\ll 1$.
+Figure 1.3: Shrinkage of principal components in the linear prediction of ridge when $\lambda=1/2$.
+The ridgeless estimator can also be expressed as a penalized LSE.
+Proposition 1.1.2 (Penalized expression of ridgeless).
+The ridgeless estimator is the only solution to the least squares problem (1.2) that is in $\operatorname{Range}(\bm{X}^{\prime})$, and it can be expressed as:
+$\hat{\bm{\theta}}_{n}^{rl}=\operatorname*{argmin}_{\bm{\theta}\in\mathbb{R}^{p}}\frac{1}{2}\left\lVert\bm{y}-\bm{X}\bm{\theta}\right\rVert_{2}^{2}+\frac{1}{2}\bm{\theta}^{\prime}\bm{\operatorname{P}}_{\operatorname{Ker}(\bm{X})}\bm{\theta}.$ (1.13)
+Proof.
+From Theorem 1.1.1, the solution set of the least squares problem (1.2) is $\hat{\bm{\theta}}_{n}^{rl}+\operatorname{Ker}(\bm{X})$, where $\hat{\bm{\theta}}_{n}^{rl}=\bm{X}^{+}\bm{y}$ is in $\operatorname{Range}(\bm{X}^{\prime})$.
+Since $\operatorname{Ker}(\bm{X})\perp\operatorname{Range}(\bm{X}^{\prime})$, $\hat{\bm{\theta}}_{n}^{rl}$ is the only least squares solution in $\operatorname{Range}(\bm{X}^{\prime})$.
+Moreover, penalty
+$h:\bm{\theta}\mapsto\bm{\theta}^{\prime}\bm{\operatorname{P}}_{\operatorname{Ker}(\bm{X})}\bm{\theta}=\bm{\theta}^{\prime}(I-\bm{X}^{+}\bm{X})\bm{\theta}$
+is zero in $\hat{\bm{\theta}}_{n}^{rl}=\bm{X}^{+}\bm{y}$, and strictly positive at any other least squares solution. We conclude that $\hat{\bm{\theta}}_{n}^{rl}$ minimizes (1.13). ∎
+Following linear transformations relate the ridgeless and the ridge estimators.
+Proposition 1.1.3 (Links between ridgeless and ridge).
+Following relations between the ridge and the ridgeless estimators hold:
+$\hat{\bm{\theta}}_{n}^{r}(\lambda)=(\bm{X}^{\prime}\bm{X}+\lambda\bm{I})^{-1}\bm{X}^{\prime}\bm{X}\hat{\bm{\theta}}_{n}^{rl},$ (1.14)
+$\hat{\bm{\theta}}_{n}^{rl}=(\bm{X}^{\prime}\bm{X})^{+}(\bm{X}^{\prime}\bm{X}+\lambda\bm{I})\hat{\bm{\theta}}_{n}^{r}(\lambda),$ (1.15)
+$\lim_{\lambda\to 0}\hat{\bm{\theta}}_{n}^{r}(\lambda)=\hat{\bm{\theta}}_{n}^{rl}.$ (1.16)
+Proof.
+Using $\bm{X}=\bm{\operatorname{P}}_{\operatorname{Range}(\bm{X})}\bm{X}$ which implies $\bm{X}^{\prime}=\bm{X}^{\prime}\bm{\operatorname{P}}_{\operatorname{Range}(\bm{X})}$, we have
+$(\bm{X}^{\prime}\bm{X}+\lambda\bm{I})^{-1}\bm{X}^{\prime}\bm{y}=(\bm{X}^{\prime}\bm{X}+\lambda\bm{I})^{-1}\bm{X}^{\prime}\bm{X}\bm{X}^{+}\bm{y},$
+and thus
+$\hat{\bm{\theta}}_{n}^{r}(\lambda)=(\bm{X}^{\prime}\bm{X}+\lambda\bm{I})^{-1}\bm{X}^{\prime}\bm{X}\hat{\bm{\theta}}_{n}^{rl}.$
+Using $\bm{X}^{+}=\bm{X}^{+}(\bm{X}^{+})^{\prime}\bm{X}^{\prime}$, we have
+$\bm{X}^{+}\bm{y}=\bm{X}^{+}(\bm{X}^{+})^{\prime}(\bm{X}^{\prime}\bm{X}+\lambda\bm{I})(\bm{X}^{\prime}\bm{X}+\lambda\bm{I})^{-1}\bm{X}^{\prime}\bm{y}.$
+Moreover, $\bm{X}^{+}({\bm{X}^{+}})^{\prime}=(\bm{X}^{\prime}\bm{X})^{+}$ implies
+$\hat{\bm{\theta}}_{n}^{rl}=(\bm{X}^{\prime}\bm{X})^{+}(\bm{X}^{\prime}\bm{X}+\lambda\bm{I})\hat{\bm{\theta}}_{n}^{r}(\lambda).$
+Finally, since $\bm{X}^{+}=\lim_{\lambda\to 0}(\bm{X}^{\prime}\bm{X}+\lambda\bm{I})^{-1}\bm{X}^{\prime}$, we have $\lim_{\lambda\to 0}\hat{\bm{\theta}}_{n}^{r}(\lambda)=\hat{\bm{\theta}}_{n}^{rl}$. ∎
+Expression (1.16) explains why estimator (1.3) is called the ridgeless estimator. The ridge and lasso estimators can be expressed as constrained least squares problems.
+Proposition 1.1.4 (Equivalence between penalized and constrained least squares).
+For $c\geq 0$, $\lambda\geq 0$, and some norm $\left\lVert\cdot\right\rVert:\mathbb{R}^{p}\to\mathbb{R}$, define:
+$\mathcal{C}(c):=\operatorname*{argmin}_{\bm{\theta}\in\mathbb{R}^{p}}\left\{\left\lVert\bm{y}-\bm{X}\bm{\theta}\right\rVert_{2}^{2}/2\ :\ \left\lVert\bm{\theta}\right\rVert\leq c\right\};$
+$\mathcal{P}(\lambda):=\operatorname*{argmin}_{\bm{\theta}\in\mathbb{R}^{p}}\left\{\left\lVert\bm{y}-\bm{X}\bm{\theta}\right\rVert_{2}^{2}/2+\lambda\left\lVert\bm{\theta}\right\rVert\right\}.$
+Then, for a given $c>0$, there exists $\lambda_{0}\geq 0$ such that $\mathcal{C}(c)\subset\mathcal{P}(\lambda_{0})$.
+Conversely, for a given $\lambda>0$, there exists $c_{0}\geq 0$ such that $\mathcal{P}(\lambda)\subset\mathcal{C}(c_{0})$.
+Proof.
+The objective function $h:\bm{\theta}\mapsto\left\lVert\bm{y}-\bm{X}\bm{\theta}\right\rVert_{2}^{2}/2$ is convex and continuous, and the constraint set $\{\bm{\theta}\in\mathbb{R}^{p}\ :\ \left\lVert\bm{\theta}\right\rVert\leq c\}$ is not empty, closed, bounded and convex.
+Therefore, a solution exists, i.e., $\mathcal{C}(c)$ is not empty; see Proposition LABEL:app:prop:existence_minimizers.
+By the KKT theorem for convex problems, $\hat{\bm{\theta}}\in\mathcal{C}(c)$ for any $c>0$ if and only if $\hat{\bm{\theta}}$ satisfies the KKT conditions, for some corresponding $\lambda_{0}\geq 0$:
+$\bm{0}\in\lambda_{0}\partial\lVert\hat{\bm{\theta}}\rVert+\bm{X}^{\prime}\bm{X}\hat{\bm{\theta}}-\bm{X}^{\prime}\bm{y},$
+$\lVert\hat{\bm{\theta}}\rVert\leq c,$
+$\lambda_{0}(\lVert\hat{\bm{\theta}}\rVert-c)=0.$
+By Theorem LABEL:thm:_fermat's_rule, the first of these conditions implies that $\hat{\bm{\theta}}\in\mathcal{P}(\lambda_{0})$.
+Now fix a $\lambda>0$ and notice that $\mathcal{P}(\lambda)$ is not empty, given that its objective function is convex, continuous and coercive; see Proposition LABEL:app:prop:existence_minimizers.
+We can thus take some $\hat{\bm{\theta}}\in\mathcal{P}(\lambda)$.
+Then, $\hat{\bm{\theta}}$ satisfies the KKT conditions of the constrained problem for $c_{0}=\lVert\hat{\bm{\theta}}\rVert$, which implies $\hat{\bm{\theta}}\in\mathcal{C}(c_{0})$.
+∎
+Note that the link between the penalty parameter $\lambda$ and the constraint parameter $c$ is not explicit.
+1.1.3 Geometric interpretation
+We illustrate the geometry of the least squares, ridge, and lasso solutions through a simple example.
+Consider the linear model (1.1), with $p=2$, $\varepsilon_{0i}\sim iiN(0,1)$, $\bm{\theta}_{0}=[1.5,0.5]^{\prime}$, $\mathbb{E}[\bm{x}_{i}\varepsilon_{0}]=\bm{0}$, and
 […]
+Figure 1.4 shows the level curves of the least squares loss function $f(\bm{\theta}):=\left\lVert\bm{y}-\bm{X}\bm{\theta}\right\rVert_{2}^{2}/2$, corresponding to values $f_{1}<f_{2}<f_{3}<f_{4}$.
+Its minimizer, or least squares solution $\hat{\bm{\theta}}_{n}^{\text{ls}}$, which coincides with the ridgeless solution $\hat{\bm{\theta}}_{n}^{rl}$, is highlighted in the figure.
+Figure 1.4: Geometry of the least squares solution.
+To illustrate the geometry of the ridge solution, we consider the constrained formulation of the ridge problem; see Proposition 1.1.4.
+Figure 1.5 demonstrates the impact of imposing the ridge constraint, represented by the sphere $\{\bm{\theta}\in\mathbb{R}^{2}\ :\ \left\lVert\bm{\theta}\right\rVert_{2}\leq c\}$ with $c=0.5$, on the least squares problem.
+The ridge solution $\hat{\bm{\theta}}_{n}^{r}$ is located at the intersection between the ridge constraint and the lower level set of the least squares loss at the lowest height (see Appendix LABEL:app_sec:convex_analysis Definition LABEL:app:def:lower_level_set) for which the intersection is non-empty.
+If the ridgeless solution $\hat{\bm{\theta}}_{n}^{rl}$ lies within the constraint boundary, then $\hat{\bm{\theta}}_{n}^{r}$ coincides with $\hat{\bm{\theta}}_{n}^{rl}$.
+Otherwise, the ridge solution $\hat{\bm{\theta}}_{n}^{r}$, by construction, is closer to the origin than $\hat{\bm{\theta}}_{n}^{rl}$, demonstrating the shrinkage effect of the ridge penalty.
+In general, $\hat{\bm{\theta}}_{n}^{r}$ is dense (i.e., contains no zero elements) with probability one.
+Figure 1.5: Geometry of the ridge solution.
+Figure 1.6 illustrates the effect of the lasso constraint, represented by the rotated square $\{\bm{\theta}\in\mathbb{R}^{2}\ :\ \left\lVert\bm{\theta}\right\rVert_{1}\leq c\}$ with $c=0.5$, on the least squares solution.
+Like the ridge solution, the lasso solution $\hat{\bm{\theta}}_{n}^{l}$ is located at the intersection of the lasso constraint and the lower level set of the least squares loss at the lowest height for which the intersection is non-empty.
+For small values of $c$, this intersection is more likely to occur along one of the coordinate axes.
+As a result, the lasso solution tends to be sparse, meaning that some components of $\hat{\bm{\theta}}_{n}^{l}$ are exactly zero.
+Figure 1.6: Geometry of the lasso solution.
+As discussed in Section 1.1, the lasso estimator serves as an approximation to the $l_{0}$-estimator (1.6).
+This relationship becomes evident through visual comparison of Figure 1.6 and Figure 1.7.
+The lasso constraint set $\{\bm{\theta}\ :\ \left\lVert\bm{\theta}\right\rVert_{1}\leq c\}$ is the convex hull (i.e., the smallest convex superset) of the constraint set underlying the $l_{0}$-estimator, which is given by: $\{\bm{\theta}\ :\ \left\lVert\bm{\theta}\right\rVert_{0}\leq c,\ \left\lVert\bm{\theta}\right\rVert_{\infty}\leq 1\}$.
+Further details on this approximation can be found in Argyriou et al.
+(2012).
+Figure 1.7: Geometry of the $l_{0}$ solution.
+To illustrate the geometry of the ridgeless solution, consider the linear model (1.1) with $p=2$, $\varepsilon_{0i}\sim\text{iid}N(0,1)$, $x_{i1}\sim\text{iid}N(0,1)$, $\mathbb{E}[x_{i1}\varepsilon_{0i}]=0$, and $x_{i2}=2x_{i1}$.
+In this case, the predictors are linearly dependent.
+As a result, the second-moment matrix of the predictors is reduced-rank:
+$\mathbb{E}[\bm{x}_{i}\bm{x}_{i}^{\prime}]=\begin{bmatrix}1&2\\ 2&4\end{bmatrix}$,
+with $\text{Rank}(\mathbb{E}[\bm{x}_{i}\bm{x}_{i}^{\prime}])=1$.
+Let $\mathbb{E}[x_{i1}y_{i}]=1$.
+The identifying condition $\mathbb{E}[\bm{x}_{i}\varepsilon_{0i}]=\bm{0}$ holds if and only if the population coefficient $\bm{\theta}_{0}$ satisfies
+$\mathbb{E}[\bm{x}_{i}\bm{x}_{i}^{\prime}]\bm{\theta}_{0}=\mathbb{E}[\bm{x}_{i}y_{i}]\quad\Longleftrightarrow\quad\begin{bmatrix}1&2\\ 2&4\end{bmatrix}\bm{\theta}_{0}=\begin{bmatrix}1\\ 2\end{bmatrix}$.
+Thus, any coefficient in the set $\bm{\theta}_{0}^{\text{rl}}+\text{Kernel}(\mathbb{E}[\bm{x}_{i}\bm{x}_{i}^{\prime}])$ satisfies this condition, where
+$\bm{\theta}_{0}^{rl}:=\mathbb{E}[\bm{x}_{i}\bm{x}_{i}^{\prime}]^{+}\mathbb{E}[\bm{x}_{i}y_{i}]=[0.2,0.4]^{\prime}.$
+If the sample size $n>\operatorname{Rank}(\mathbb{E}[\bm{x}_{i}\bm{x}_{i}^{\prime}])$, then $\operatorname{Ker}(\bm{X})\supset\operatorname{Ker}(\mathbb{E}[\bm{x}_{i}\bm{x}_{i}^{\prime}])$, and the same issue arises in the finite-sample least squares problem, where the objective function $f(\bm{\theta})$ is minimized at any point on the affine set
+$\hat{\bm{\theta}}_{n}^{rl}+\operatorname{Ker}(\bm{X}).$
+Figure 1.8 depicts the level curves of $f(\bm{\theta})$ at $f(\hat{\bm{\theta}}_{n}^{\text{rl}})=f_{1}<f_{2}<f_{3}$.
+These curves are parallel lines, unlike the typical ellipses seen in full-rank cases.
+The ridgeless estimator is the minimum $l_{2}$-norm solution to the least squares problem, as expected from its construction.
+Figure 1.8: Geometry of the ridgeless solution.
+1.1.4 Computation of lasso
+By Theorem LABEL:thm:_fermat's_rule, a function $f:\mathbb{R}^{p}\to\mathbb{R}$ is minimized at $\bm{\theta}^{*}\in\mathbb{R}^{p}$ if and only if $\bm{0}\in\partial f(\bm{\theta}^{*})$.
+In the lasso problem, the objective function
+$f^{l}:\mathbb{R}^{p}\to\mathbb{R};\bm{\theta}\mapsto\frac{1}{2n}\left\lVert\bm{y}-\bm{X}\bm{\theta}\right\rVert_{2}^{2}+\lambda\left\lVert\bm{\theta}\right\rVert_{1}$ (1.17)
+contains the $l_{1}$-norm, which makes $f^{l}$ non-smooth.
+[^5] Notice that $f^{l}$ is the objective function in (1.5), multiplied by $1/n$.
+This term does not show up in the penalization term as it is absorbed by $\lambda$.
+As a result, the subdifferential of $f^{l}$ at a minimizer $\hat{\bm{\theta}}_{n}^{l}(\lambda)$ is not necessarily a singleton, implying that the lasso estimator may not be unique.
+Moreover, due to the complexity of $\partial f^{l}(\hat{\bm{\theta}}_{n}^{l}(\lambda))$, no closed-form solution exists in general for the lasso estimator.
+Fortunately, Proposition 1.1.4 implies that the lasso problem is a quadratic program with a convex constraint, which allows for the computation of the lasso estimator using various quadratic programming algorithms.
+One particularly simple and effective method is the cyclical coordinate descent algorithm, which minimizes the convex objective function by iterating through each coordinate independently.
+This approach provides insight into how the lasso solution is obtained.
+Consider the soft-thresholding operator for a given $\lambda>0$, which is defined as the function
+$\mathcal{S}_{\lambda}:\mathbb{R}\to\mathbb{R};\eta\mapsto\begin{cases}\eta-\lambda,&\eta>\lambda\\ 0,&\eta\in[-\lambda,\lambda]\\ \eta+\lambda,&\eta<-\lambda\end{cases}.$
+This operator is illustrated in Figure 1.9.
+Figure 1.9: Soft-thresholding operator.
+The soft-thresholding operator provides a direct way to compute the lasso estimator in a univariate regression model, i.e., when there is only one predictor.
+Proposition 1.1.5 (Lasso solution for univariate regression).
+Given $\lambda>0$ and $\bm{X}_{1}\in\mathbb{R}^{n}$ such that $\bm{X}_{1}^{\prime}\bm{X}_{1}>0$, we have
+$\hat{\theta}_{n}^{l}(\lambda):=\operatorname*{argmin}_{\theta\in\mathbb{R}}\left\{\frac{1}{2n}\left\lVert\bm{y}-\bm{X}_{1}\theta\right\rVert_{2}^{2}+\lambda|\theta|\right\}=\frac{\mathcal{S}_{\lambda}(\bm{X}_{1}^{\prime}\bm{y}/n)}{\bm{X}_{1}^{\prime}\bm{X}_{1}/n}.$
+Proof.
+The subdifferential of $f^{l}:\theta\mapsto\frac{1}{2n}\left\lVert\bm{y}-\bm{X}_{1}\theta\right\rVert_{2}^{2}+\lambda|\theta|$ at $\hat{\theta}\in\mathbb{R}$ reads
+[…]
+where $a:=\bm{X}_{1}^{\prime}\bm{y}/n$ and $b:=\bm{X}_{1}^{\prime}\bm{X}_{1}/n$.
+From Theorem LABEL:thm:_fermat's_rule, and the subdifferential of the absolute value function (Appendix LABEL:app_sec:convex_analysis, Example LABEL:subdifferetnial_absolute_value), $\hat{\theta}^{l}$ is a minimizer of $f^{l}$ if and only if
+$0\in\partial f^{l}(\hat{\theta}^{l})\quad\Longleftrightarrow\quad a\in b\hat{\theta}^{l}+\begin{cases}\{\lambda\},&\hat{\theta}^{l}>0\\ [-\lambda,\lambda],&\hat{\theta}^{l}=0\\ \{-\lambda\},&\hat{\theta}^{l}<0\end{cases}.$
+This condition reads: (i) if $\hat{\theta}^{l}>0$, then $\hat{\theta}^{l}=(a-\lambda)/b$, implying $a>\lambda$; (ii) if $\hat{\theta}^{l}=0$, then $\lambda\leq a\leq\lambda$; and (iii) if $\hat{\theta}^{l}<0$, then $\hat{\theta}^{l}=(a+\lambda)/b$, implying $a<-\lambda$.
+These cases are summarized by $\hat{\theta}^{l}=\mathcal{S}_{\lambda}(a)/b$.
+∎
+Proposition 1.1.5 can be used to show that the $j$-th coordinate of the lasso solution in a multivariate regression model, i.e., when there is more than just one predictor, satisfies an expression based on the soft-thresholding operator applied to the residual of a lasso regression of $\bm{y}$ onto the predictors $\bm{X}_{k}$ at position $k\neq j$.
+Theorem 1.1.4 (Lasso solution).
+Let $\bm{X}_{j}$ denote the $j$-th column of $\bm{X}$ and $\bm{X}_{(-j)}$ denote $\bm{X}$ without the $j$-th column.
+Assume that $\bm{X}_{j}^{\prime}\bm{X}_{j}>0$ for all $j=1,\ldots,p$.
+Then, given $\lambda>0$, any lasso solution $\hat{\bm{\theta}}_{n}^{l}(\lambda)$ is such that for all $j=1,\ldots,p$:
+$\hat{\theta}_{n,j}^{l}(\lambda)=\operatorname*{argmin}_{\theta\in\mathbb{R}}\left\{\frac{1}{2n}\left\lVert\bm{e}_{j}-\bm{X}_{j}\theta\right\rVert_{2}^{2}+\lambda|\theta|\right\}=\frac{\mathcal{S}_{\lambda}(\bm{X}_{j}^{\prime}\bm{e}_{j}/n)}{\bm{X}_{j}^{\prime}\bm{X}_{j}/n},$ (1.18)
+where $\hat{\theta}_{n,j}^{l}(\lambda)$ is the $j$-th element of $\hat{\bm{\theta}}_{n}^{l}(\lambda)$, $\hat{\bm{\theta}}_{n,(-j)}^{l}(\lambda)$ is $\hat{\bm{\theta}}_{n}^{l}(\lambda)$ without the $j$-th element, and $\bm{e}_{j}:=\bm{y}-\bm{X}_{(-j)}\hat{\bm{\theta}}_{n,(-j)}^{l}(\lambda)$.
+Proof.
+The subdifferential of the lasso objective function $f^{l}$ defined in (1.17) at $\hat{\bm{\theta}}\in\mathbb{R}^{p}$ is
+$\partial f^{l}(\hat{\bm{\theta}})=(\bm{X}^{\prime}\bm{X}/n)\hat{\bm{\theta}}-\bm{X}^{\prime}\bm{y}/n+\lambda\partial\lVert\hat{\bm{\theta}}\rVert_{1},$
+where
+$\partial\lVert\hat{\bm{\theta}}\rVert_{1}=\{\bm{v}\in\mathbb{R}^{p}\ :\ v_{j}\in\partial|\hat{\theta}_{j}|\ \text{for all}\ j=1,\ldots,p\}.$
+and the subdifferential of the absolute value function $|\cdot|$ is given in Appendix LABEL:app_sec:convex_analysis, Example LABEL:subdifferetnial_absolute_value.
+From Theorem LABEL:thm:_fermat's_rule, a minimizer $\hat{\bm{\theta}}_{n}^{l}(\lambda)$ of $f^{l}$ satisfies
+$\bm{0}\in\partial f^{l}(\hat{\bm{\theta}}_{n}^{l}(\lambda))\quad\Longleftrightarrow\quad\bm{X}^{\prime}\bm{y}/n\in\bm{X}^{\prime}\bm{X}/n\hat{\bm{\theta}}_{n}^{l}(\lambda)+\lambda\partial\lVert\hat{\bm{\theta}}_{n}^{l}(\lambda)\rVert_{1}.$
+This condition holds if and only if for all $j=1,\ldots,p$:
+$\bm{X}_{j}^{\prime}\bm{y}/n$ $\in\bm{X}_{j}^{\prime}\bm{X}/n\hat{\bm{\theta}}_{n}^{l}(\lambda)+\lambda\partial\lvert\hat{\theta}_{n,j}^{l}(\lambda)\rvert\quad\Longleftrightarrow$
+$\bm{X}_{j}^{\prime}\bm{e}_{j}/n$ $\in\bm{X}_{j}^{\prime}\bm{X}_{j}/n\hat{\theta}_{n,j}^{l}(\lambda)+\lambda\partial\lvert\hat{\theta}_{n,j}^{l}(\lambda)\rvert\quad\Longleftrightarrow$
+$\hat{\theta}_{n,j}^{l}(\lambda)$ $=\frac{\mathcal{S}_{\lambda}(\bm{X}_{j}^{\prime}\bm{e}_{j}/n)}{\bm{X}_{j}^{\prime}\bm{X}_{j}/n},$
+where the first double implication follows from
+$\bm{X}_{j}^{\prime}\bm{X}\hat{\bm{\theta}}_{n}^{l}(\lambda)=\sum_{k=1}^{p}\bm{X}_{j}^{\prime}\bm{X}_{k}\hat{\theta}_{n,k}^{l}(\lambda)=\bm{X}_{j}^{\prime}\bm{X}_{(-j)}\hat{\bm{\theta}}_{n,(-j)}^{l}(\lambda)+\bm{X}_{j}^{\prime}\bm{X}_{j}\hat{\theta}_{n,j}^{l}(\lambda),$
+and the last double implication follows from Proposition 1.1.5 since, by Theorem LABEL:thm:_fermat's_rule,
+$\bm{X}_{j}^{\prime}\bm{e}_{j}/n$ $\in\bm{X}_{j}^{\prime}\bm{X}_{j}/n\hat{\theta}_{n,j}^{l}(\lambda)+\lambda\partial\lvert\hat{\theta}_{n,j}^{l}(\lambda)\rvert\quad\Longleftrightarrow$
+$\hat{\theta}_{n,j}^{l}(\lambda)$ $=\operatorname*{argmin}_{\theta\in\mathbb{R}}\left\{\frac{1}{2n}\left\lVert\bm{e}_{j}-\bm{X}_{j}\theta\right\rVert_{2}^{2}+\lambda|\theta|\right\}.$
+Theorem 1.1.4 suggests that the lasso solution can be computed by a cyclical coordinate minimization algorithm.
+This method is an iterative algorithm that, given a candidate solution $\hat{\bm{\theta}}^{(t)}$ at iteration $t$, updates a single coordinate $j$ as
+$\hat{\theta}_{j}^{(t+1)}=\operatorname*{argmin}_{\theta\in\mathbb{R}}f(\hat{\theta}_{1}^{(t)},\ldots,\hat{\theta}_{j-1}^{(t)},\theta,\hat{\theta}_{j+1}^{(t)},\ldots,\hat{\theta}_{p}^{(t)}),$
+while keeping all other coordinates fixed, i.e., $\hat{\theta}_{k}^{(t+1)}=\hat{\theta}_{k}^{(t)}$ for $k\neq j$.
+A typical choice for the lasso solution would be to cycle through the coordinates in their natural order: from $1$ to $p$.
+The coordinate descent algorithm is guaranteed to converge to a global minimizer of any convex cost function $f:\mathbb{R}^{p}\to\mathbb{R}$ satisfying the additive decomposition:
+$f:\bm{\theta}\mapsto g(\bm{\theta})+\sum_{j=1}^{p}h_{j}(\theta_{j}),$
+where $g:\mathbb{R}^{p}\to\mathbb{R}$ is differentiable and convex, and the univariate functions $h_{j}:\mathbb{R}^{p}\to\mathbb{R}$ are convex (but not necessarily differentiable); see Tseng (2001).
+What makes this algorithm work for the lasso problem is the fact that objective function (1.17) satisfies this separable structure.
+Remark 5.
+If the predictors are measured in different units, it is recommended to standardize them so that $\bm{X}_{j}^{\prime}\bm{X}_{j}=1$ for all $j$.
+In this case, the lasso update (1.18) has the simpler form:
+$\hat{\theta}_{n,j}^{l}(\lambda)=\mathcal{S}_{\lambda}(\bm{X}_{j}^{\prime}\bm{e}_{j}/n).$
+Algorithm 1 summarizes the pseudo-code of the cyclical coordinate descent algorithm for computing the lasso estimator.
+This algorithm proceeds by cyclically applying the soft-thresholding update in (1.18) for each coordinate, simultaneously updating the residuals $\bm{e}_{j}:=\bm{y}-\bm{X}_{(-j)}\hat{\bm{\theta}}_{n,(-j)}^{l}(\lambda)$.
+The ridgeless or the ridge estimators can be used to initialize the procedure.
+Algorithm 1 Cyclical coordinate descent method for the lasso estimator.
+$\bm{y}\in\mathbb{R}^{n}$, $\bm{X}\in\mathbb{R}^{n\times p}$ with $\bm{X}_{j}^{\prime}\bm{X}_{j}>0$ for all $j=1,\ldots,p$
+Penalty parameter $\lambda>0$
+Initial estimator $\hat{\bm{\theta}}_{n}^{i}$ (e.g., ridgeless or ridge)
+Maximum number of iterations $T$ and tolerance $\varepsilon>0$
+Standardize: $\bm{y}^{\prime}\bm{1}=0$, $\bm{X}^{\prime}\bm{1}=\bm{0}$, and $\bm{X}_{j}^{\prime}\bm{X}_{j}/n=1$ for all $j$
+$\hat{\bm{\theta}}^{(0)}\leftarrow\hat{\bm{\theta}}_{n}^{i}$
+for $t=1,\ldots,T$ do
+  for $j=1,\ldots,p$ do
+   $\bm{e}_{j}\leftarrow\bm{y}-\bm{X}_{(-j)}\,\hat{\bm{\theta}}_{(-j)}^{(t-1)}$
+   $\hat{\theta}_{j}^{(t)}\leftarrow\mathcal{S}_{\lambda}\!\left(\bm{X}_{j}^{\prime}\bm{e}_{j}/n\right)$
+  end for
+  if $\|\hat{\bm{\theta}}^{(t)}-\hat{\bm{\theta}}^{(t-1)}\|_{\infty}<\varepsilon$ then
+   break
+  end if
+end for
+Output: $\hat{\bm{\theta}}^{(t)}$
+In practice, it is often desirable to compute the lasso solution not for a single fixed value of $\lambda$, but for the entire solution path over a range of $\lambda$ values.
+A common approach begins by selecting a value of $\lambda$ just large enough that the only optimal solution is the zero vector.
+This value is denoted as $\lambda_{\max}=\max_{j}\{|\bm{X}_{j}^{\prime}\bm{y}|/n\}$.
+Starting from $\lambda_{\max}$, we progressively decrease $\lambda$ and, at each step, run coordinate descent to convergence, initializing the algorithm with the solution obtained for the previous value of $\lambda$ (a warm start).
+In this way, we can efficiently compute the solutions over a grid of $\lambda$ values.
+This approach is known as pathwise coordinate descent.
+Coordinate descent is particularly efficient for the lasso because the update rule (1.18) is available in closed form, eliminating the need for iterative searches along each coordinate.
+Additionally, the algorithm exploits the inherent sparsity of the problem: for sufficiently large values of $\lambda$, most coefficients will be zero and will remain unchanged.
+There are also computational strategies that can predict the active set of variables, significantly speeding up the algorithm.
+More details on the pathwise coordinate descent algorithm for lasso can be found in Friedman et al.
+(2007).
+Homotopy methods are another class of techniques for solving the lasso estimator.
+They produce the entire path of solutions in a sequential fashion, starting at zero.
+A homotopy method that is particularly efficient at computing the entire lasso path is the least angle regression (LARS) algorithm; see Efron et al.
+(2004).
+1.1.5 Finite-sample properties of ridgeless and ridge
+This section presents finite-sample expressions and bounds for the bias, variance, MSE and MPR of the LSE, ridgeless and ridge estimators.
+The main underlying assumption is that linear model $y=\bm{x}^{\prime}\bm{\theta}_{0}+\varepsilon_{0}$ satisfies the typical regression condition $\mathbb{E}[\bm{x}\varepsilon_{0}]=\bm{0}$.
+Furthermore, we work with a fixed design matrix $\bm{X}$ (or equivalently we work conditionally on $\bm{X}$).
+The next proposition derives the bias, MSE and MPR of the LSE when it is well-defined, that is, when $\operatorname{Rank}(\bm{X})=p$, which implies $p\leq n$.
+Proposition 1.1.6 (Finite-sample properties of LSE (fixed design)).
+Assume that the linear model $y=\bm{x}^{\prime}\bm{\theta}_{0}+\varepsilon_{0}$ holds with $\mathbb{E}[\varepsilon_{0}\mid\bm{x}]=\bm{0}$.
+Fix a realization $\bm{X}\in R^{n\times p}$ with $\operatorname{Rank}(\bm{X})=p$.
+Then (throughout, expectations are understood conditional on the fixed design $X$.):
+  1. (i)
+    The LSE is unbiased: $\mathbb{E}[\hat{\bm{\theta}}_{n}^{ls}]=\bm{\theta}_{0}.$
+  2. (ii)
+    The variance of the LSE is given by
+    $\operatorname{\mathbb{V}ar}[\hat{\bm{\theta}}_{n}^{ls}]=(\bm{X}^{\prime}\bm{X})^{-1}\bm{X}^{\prime}\operatorname{\mathbb{V}ar}[\bm{\varepsilon}_{0}]\bm{X}(\bm{X}^{\prime}\bm{X})^{-1}.$
+Further let $\operatorname{\mathbb{V}ar}[\bm{\varepsilon}_{0}]=\sigma^{2}\bm{I}$ with $\sigma>0$. Then:
+  1. (iii)
+    $\operatorname{\mathbb{V}ar}[\hat{\bm{\theta}}_{n}^{ls}]=\sigma^{2}(\bm{X}^{\prime}\bm{X})^{-1}$.
+  2. (iv)
+    The LSE is the best linear unbiased estimator, in the sense that $\operatorname{\mathbb{V}ar}[\tilde{\bm{\theta}}_{n}]-\operatorname{\mathbb{V}ar}[\hat{\bm{\theta}}_{n}^{ls}]$ is positive semi-definite for any other unbiased linear estimator $\tilde{\bm{\theta}}_{n}$, i.e., $\tilde{\bm{\theta}}_{n}=\bm{A}\bm{y}$ for some $\bm{A}\in\mathbb{R}^{p\times n}$.
+  3. (v)
+    The MSE of the LSE is given by: $\operatorname{MSE}(\hat{\bm{\theta}}_{n}^{ls},\bm{\theta}_{0})=\frac{\sigma^{2}}{n}\sum_{j=1}^{p}\frac{1}{\lambda_{j}}$, where $\lambda_{1}\geq\ldots\geq\lambda_{p}>0$ are the eigenvalues of $\bm{X}^{\prime}\bm{X}/n$.
+Therefore,
+    $\operatorname{MSE}(\hat{\bm{\theta}}_{n}^{ls},\bm{\theta}_{0})\leq\frac{\sigma^{2}p}{\lambda_{p}n}.$
+  4. (vi)
+    The mean predictive risk of the LSE is given by:
+    $\operatorname{MPR}(\hat{\bm{\theta}}_{n}^{ls},\bm{\theta}_{0})=p\sigma^{2}/n.$
+Proof.
+  1. (i)
+    From the closed form expression (1.8):
+    $\hat{\bm{\theta}}_{n}^{ls}=(\bm{X}^{\prime}\bm{X})^{-1}\bm{X}^{\prime}\bm{\varepsilon}_{0}+\bm{\theta}_{0}.$
+    Thus, unbiasedness follows directly since:
+    $\mathbb{E}[\hat{\bm{\theta}}_{n}^{ls}]=\bm{\theta}_{0}+(\bm{X}^{\prime}\bm{X})^{-1}\mathbb{E}[\bm{X}^{\prime}\bm{\varepsilon}_{0}]=\bm{\theta}_{0}.$
+  2. (ii)
+    The closed form expression of $\hat{\bm{\theta}}_{n}^{ls}$ immediately implies the expression
+    $\operatorname{\mathbb{V}ar}[\hat{\bm{\theta}}_{n}^{ls}]=$ $\operatorname{\mathbb{V}ar}[(\bm{X}^{\prime}\bm{X})^{-1}\bm{X}^{\prime}\bm{\varepsilon}_{0}]$
+    $=$ $(\bm{X}^{\prime}\bm{X})^{-1}\bm{X}^{\prime}\operatorname{\mathbb{V}ar}[\bm{\varepsilon}_{0}]\bm{X}(\bm{X}^{\prime}\bm{X})^{-1}.$
+  3. (iii)
+    Das ist trivial.
+  4. (iv)
+    A linear estimator $\tilde{\bm{\theta}}_{n}=\bm{A}\bm{y}$ is unbiased if and only if $\bm{A}\bm{X}=\bm{I}$.
+Let $\bm{M}:=\bm{X}(\bm{X}^{\prime}\bm{X})^{-1}\bm{X}^{\prime}$, and notice that $(\bm{I}-\bm{M})(\bm{I}-\bm{M})=(\bm{I}-\bm{M})$, i.e., $(\bm{I}-\bm{M})$ is idempotent.
+It follows that:
+    $\operatorname{\mathbb{V}ar}[\tilde{\bm{\theta}}_{n}]-\operatorname{\mathbb{V}ar}[\hat{\bm{\theta}}_{n}^{ls}]=$ $\sigma^{2}(\bm{A}\bm{A}^{\prime}-(\bm{X}^{\prime}\bm{X})^{-1})$
+    $=$ $\sigma^{2}(\bm{A}\bm{A}^{\prime}-\bm{A}\bm{X}(\bm{X}^{\prime}\bm{X})^{-1}\bm{X}^{\prime}\bm{A}^{\prime})$
+    $=$ $\sigma^{2}\bm{A}(\bm{I}-\bm{M})\bm{A}^{\prime}$
+    $=$ $\sigma^{2}[\bm{A}(\bm{I}-\bm{M})][\bm{A}(\bm{I}-\bm{M})]^{\prime},$
+    which is positive semi definite.
+  5. (v)
+    Using the linearity of the $\operatorname{Trace}$ operator and the SVD decomposition of $\bm{X}$ in Definition 18:
+    $\mathbb{E}[\lVert\hat{\bm{\theta}}_{n}^{ls}-\bm{\theta}_{0}\rVert_{2}^{2}]=$ $\mathbb{E}[\operatorname{Trace}((\hat{\bm{\theta}}_{n}^{ls}-\bm{\theta}_{0})(\hat{\bm{\theta}}_{n}^{ls}-\bm{\theta}_{0})^{\prime})]$
+    $=$ $\operatorname{Trace}(\operatorname{\mathbb{V}ar}[\hat{\bm{\theta}}_{n}^{ls}])$
+    $=$ $\frac{\sigma^{2}}{n}\operatorname{Trace}((\bm{X}^{\prime}\bm{X}/n)^{-1})$
+    $=$ $\frac{\sigma^{2}}{n}\operatorname{Trace}(\bm{V}(\bm{S}^{\prime}\bm{S}/n)^{-1}\bm{V}^{\prime})=\frac{\sigma^{2}}{n}\sum_{j=1}^{p}\frac{1}{\lambda_{j}}.$
+  6. (vi)
+    Simple computations give
+    $\mathbb{E}[\lVert\operatorname{lm}(\bm{X},\hat{\bm{\theta}}_{n}^{ls})$ $-\operatorname{lm}(\bm{X},\bm{\theta}_{0})\rVert_{2}^{2}/n]$
+    $=$ $\mathbb{E}[\lVert\bm{X}(\hat{\bm{\theta}}_{n}^{ls}-\bm{\theta}_{0})\rVert_{2}^{2}/n]$
+    $=$ $\mathbb{E}[\operatorname{Trace}((\hat{\bm{\theta}}_{n}^{ls}-\bm{\theta}_{0})^{\prime}\bm{X}^{\prime}\bm{X}/n(\hat{\bm{\theta}}_{n}^{ls}-\bm{\theta}_{0}))]$
+    $=$ $\mathbb{E}[\operatorname{Trace}((\hat{\bm{\theta}}_{n}^{ls}-\bm{\theta}_{0})(\hat{\bm{\theta}}_{n}^{ls}-\bm{\theta}_{0})^{\prime}\bm{X}^{\prime}\bm{X}/n)]$
+    $=$ $\operatorname{Trace}(\operatorname{\mathbb{V}ar}[\hat{\bm{\theta}}_{n}^{ls}]\bm{X}^{\prime}\bm{X}/n)=\sigma^{2}p/n.$
+This proposition shows that the LSE’s accuracy decreases:
+  • as the variance $\sigma^{2}$ of the error term increases;
+  • as the number of predictors per observation $p/n$ increases;
+  • as the ”degree of singularity” of the design matrix $1/\lambda_{p}$ increases.
+Remark 6 (Why eigenvalues of $\bm{X}^{\prime}\bm{X}/n$).
+We express the results in Proposition 1.1.6 in terms of the eigenvalues $\lambda_{1}\geq\cdots\geq\lambda_{p}>0$ of the normalized Gram matrix $\bm{X}^{\prime}\bm{X}/n$, rather than those of $\bm{X}^{\prime}\bm{X}$, because the eigenvalues of $\bm{X}^{\prime}\bm{X}/n$ remain on an $O(1)$ scale as $n$ grows; see, e.g., Vershynin (2018).
+This keeps the $1/n$ convergence rate of the MSE explicit.
+In random design settings with i.i.d.
+rows $\bm{x}_{i}$ satisfying $\mathbb{E}[\|\bm{x}_{i}\|_{2}^{2}]<\infty$ and $\operatorname{\mathbb{V}ar}(\bm{x}_{i})=\bm{\Sigma}$ positive definite, the Law of Large Numbers implies $\bm{X}^{\prime}\bm{X}/n\xrightarrow{p}\bm{\Sigma}$.
+By continuity of the eigenvalue map (see, e.g., Weyl’s theorem; Horn and Johnson, 2012), it follows that $\lambda_{j}(\bm{X}^{\prime}\bm{X}/n)\xrightarrow{p}\lambda_{j}(\bm{\Sigma})$ for each $j=1,\ldots,p$.
+Hence, the eigenvalues of $\bm{X}^{\prime}\bm{X}/n$ remain $O_{p}(1)$, while those of $\bm{X}^{\prime}\bm{X}$ grow linearly with $n$.
+In fixed design analyses one typically assumes uniform spectral bounds,
+$0<\underline{\lambda}\leq\lambda_{p}\big(\bm{X}^{\prime}\bm{X}/n\big)\leq\lambda_{1}\big(\bm{X}^{\prime}\bm{X}/n\big)\leq\overline{\lambda}<\infty,$
+which plays the same role.
+We now aim to drop the requirement that $\operatorname{Rank}(\bm{X})=p$, to allow for high-dimensional settings where:
+  • $n<p$, or even
+  • $\operatorname{Rank}(\mathbb{E}[\bm{x}\bm{x}^{\prime}])=r_{0}$, not necessarily equals to $p$.
+Notice that the requirement $\operatorname{Rank}(\mathbb{E}[\bm{x}\bm{x}^{\prime}])=r_{0}$ implies $\operatorname{Rank}(\bm{X})\ \leq\ \min\{n,r_{0}\}$:
+Proposition 1.1.7 (Rank bound under low-rank second moments).
+Let $\bm{x}\in\mathbb{R}^{p}$ satisfy $\mathbb{E}\|\bm{x}\|_{2}^{2}<\infty$ and let $\bm{\Sigma}:=\mathbb{E}[\bm{x}\bm{x}^{\prime}]$ be positive semi-definite with $\operatorname{Rank}(\bm{\Sigma})=r_{0}$.
+Fix a realization $\bm{X}\in\mathbb{R}^{n\times p}$ with rows $\bm{x}_{1}^{\prime},\ldots,\bm{x}_{n}^{\prime}$, where each $\bm{x}_{i}$ has the same distribution as $\bm{x}$.
+Then, almost surely,
+$\operatorname{Rank}(\bm{X})\ \leq\ \min\{n,r_{0}\}.$
+Proof.
+Since $\mathbb{E}\|\bm{x}\|_{2}^{2}<\infty$, $\bm{\Sigma}$ is well-defined and positive semi-definite. If $\bm{a}\in\operatorname{Ker}(\bm{\Sigma})$, then
+$0=\bm{a}^{\prime}\bm{\Sigma}\bm{a}=\mathbb{E}\big[(\bm{a}^{\prime}\bm{x})^{2}\big]\quad\Longrightarrow\quad\bm{a}^{\prime}\bm{x}=0$
+almost surely.
+Hence $\bm{x}\in\operatorname{Range}(\bm{\Sigma})$ almost surely, so for each $i\in\{1,\ldots,n\}$ we also have $\bm{x}_{i}\in\operatorname{Range}(\bm{\Sigma})$ almost surely (they share the same distribution), i.e., all rows of $\bm{X}$ lie in $\operatorname{Range}(\bm{\Sigma})$ almost surely.
+Therefore,
+$\operatorname{Rank}(\bm{X})\ \leq\ \dim(\operatorname{Range}(\bm{\Sigma}))\ =\ r_{0}$
+almost surely. Combining with the trivial bound $\operatorname{Rank}(\bm{X})\leq n$ (notice that $r_{0}\leq p$ by construction) yields $\operatorname{Rank}(\bm{X})\leq\min\{n,r_{0}\}$ almost surely. ∎
+We next show that, when $\operatorname{Rank}(\mathbb{E}[\bm{x}\bm{x}^{\prime}])=r_{0}<p$, the typical linear regression condition $\mathbb{E}[\bm{x}\varepsilon_{0}]=\bm{0}$ no longer identifies a unique estimand $\bm{\theta}_{0}$.
+Proposition 1.1.8.
+Let $y\in\mathbb{R}$ and $\bm{x}\in\mathbb{R}^{p}$ be random variables with $\mathbb{E}\|\bm{x}\|_{2}^{2}<\infty$ and $\mathbb{E}[y^{2}]<\infty$.
+For any $\bm{\theta}\in\mathbb{R}^{p}$ let $\varepsilon(\bm{\theta}):=y-\bm{x}^{\prime}\bm{\theta}$.
+Consider the set
+$S:=\{\bm{\theta}_{0}\in\mathbb{R}^{p}:\ y=\bm{x}^{\prime}\bm{\theta}_{0}+\varepsilon(\bm{\theta}_{0}),\ \mathbb{E}[\bm{x}\varepsilon(\bm{\theta}_{0})]=\bm{0}\}.$
+Then $S$ is nonempty if and only if $\mathbb{E}[\bm{x}y]\in\operatorname{Range}(\mathbb{E}[\bm{x}\bm{x}^{\prime}])$, in which case
+$S=(\mathbb{E}[\bm{x}\bm{x}^{\prime}])^{+}\mathbb{E}[\bm{x}y]+\operatorname{Ker}(\bm{\Sigma}).$
+Proof.
+By definition, $y=\bm{x}^{\prime}\bm{\theta}+\varepsilon(\bm{\theta})$ holds for every $\bm{\theta}$. Thus the restriction in $S$ is the moment condition
+$\mathbb{E}[\bm{x}\varepsilon(\bm{\theta})]=\mathbb{E}[\bm{x}(y-\bm{x}^{\prime}\bm{\theta})]=\bm{0}\quad\Longleftrightarrow\quad\mathbb{E}[\bm{x}\bm{x}^{\prime}]\bm{\theta}=\mathbb{E}[\bm{x}y].$
+With $\bm{\Sigma}:=\mathbb{E}[\bm{x}\bm{x}^{\prime}]$ (symmetric positive semi-definite) and $\bm{b}:=\mathbb{E}[\bm{x}y]$ (well-defined by the moment assumptions), this is the linear system $\bm{\Sigma}\bm{\theta}=\bm{b}$.
+Such a system admits a solution if and only if $\bm{b}\in\operatorname{Range}(\bm{\Sigma})$; when this holds, the set of all solutions is
+$\{\bm{\theta}\in\mathbb{R}^{p}\ :\ \bm{\Sigma}\bm{\theta}=\bm{b}\}=\bm{\Sigma}^{+}\bm{b}+\operatorname{Ker}(\bm{\Sigma}).$
+However, when $S$ is nonempty, the unique element in $S\cap\operatorname{Range}(\mathbb{E}[\bm{x}\bm{x}^{\prime}])$ remains well defined even when $\operatorname{Rank}(\mathbb{E}[\bm{x}\bm{x}^{\prime}])<p$.
+Moreover, when $\operatorname{Rank}(\mathbb{E}[\bm{x}\bm{x}^{\prime}])=p$, this element coincides with the population parameter $\bm{\theta}_{0}$ by construction.
+We define it as follows.
+Definition 19 (Ridgeless estimand).
+The ridgeless estimand is the parameter vector
+$\bm{\theta}_{0}^{rl}:=\mathbb{E}[\bm{x}\bm{x}^{\prime}]^{+}\mathbb{E}[\bm{x}y],$
+which, by construction, satisfies $\bm{\theta}_{0}^{rl}\in\operatorname{Range}(\mathbb{E}[\bm{x}\bm{x}^{\prime}])$.
+[^6] The fact that $\bm{\theta}_{0}^{rl}\in\operatorname{Range}(\mathbb{E}[\bm{x}\bm{x}^{\prime}])$ follows from the identity $\mathbb{E}[\bm{x}\bm{x}^{\prime}]^{+}=\operatorname{P}_{\operatorname{Range}(\mathbb{E}[\bm{x}\bm{x}^{\prime}])}\mathbb{E}[\bm{x}\bm{x}^{\prime}]^{+}$.
+The ridgeless estimand is also used in Section 1.1.3.
+We can now extend Proposition 1.1.6 to the fixed design setting where $\operatorname{Rank}(\bm{X})\leq p$.
+Proposition 1.1.9 (Finite-sample properties of ridgeless (fixed design)).
+Assume that the linear model $y=\bm{x}^{\prime}\bm{\theta}_{0}+\varepsilon_{0}$ holds for some $\bm{\theta}_{0}\in\mathbb{R}^{p}$ with $\mathbb{E}[\varepsilon_{0}\mid\bm{x}]=0$, and the finite second moments $\mathbb{E}[\|\bm{x}\|_{2}^{2}]<\infty$ and $\mathbb{E}[y^{2}]<\infty$ are met.
+Denote $\bm{\Sigma}:=\mathbb{E}[\bm{x}\bm{x}^{\prime}]$, and let $r_{0}:=\operatorname{Rank}(\bm{\Sigma})$.
+Fix a realization $\bm{X}\in\mathbb{R}^{n\times p}$ with rows $\bm{x}_{1}^{\prime},\ldots,\bm{x}_{n}^{\prime}$, where each $\bm{x}_{i}$ has the same distribution as $\bm{x}$.
+Then (throughout, expectations and variances are understood conditionally on the fixed design $\bm{X}$):
+  1. (i)
+    $\mathbb{E}[\hat{\bm{\theta}}_{n}^{rl}]=\bm{P}_{\operatorname{Range}(\bm{X}^{\prime})}\bm{\theta}_{0}^{rl}$.
+If $\operatorname{Range}(\bm{X}^{\prime})=\operatorname{Range}(\bm{\Sigma})$ (which implies $n\geq r_{0}$), then the ridgeless estimator is unbiased: $\mathbb{E}[\hat{\bm{\theta}}_{n}^{rl}]=\bm{\theta}_{0}^{rl}$.
+  2. (ii)
+    $\operatorname{\mathbb{V}ar}[\hat{\bm{\theta}}_{n}^{rl}]=\bm{X}^{+}\operatorname{\mathbb{V}ar}[\bm{\varepsilon}(\bm{\theta}_{0}^{rl})](\bm{X}^{+})^{\prime}$, where $\operatorname{\mathbb{V}ar}[\bm{\varepsilon}(\bm{\theta}_{0}^{rl})]=\operatorname{\mathbb{V}ar}[\bm{\varepsilon}_{0}]$.
+Further suppose $\operatorname{\mathbb{V}ar}[\bm{\varepsilon}(\bm{\theta}_{0}^{rl})]=\sigma^{2}\bm{I}$ with $\sigma>0$, and let $r:=\operatorname{Rank}(\bm{X})$. Then:
+  1. (iii)
+    $\operatorname{\mathbb{V}ar}[\hat{\bm{\theta}}_{n}^{rl}]=\sigma^{2}(\bm{X}^{\prime}\bm{X})^{+}$.
+  2. (iv)
+    The mean squared error satisfies
+    $\operatorname{MSE}(\hat{\bm{\theta}}_{n}^{rl},\bm{\theta}_{0}^{rl})=\frac{\sigma^{2}}{n}\sum_{j=1}^{r}\frac{1}{\lambda_{j}}+\|\bm{P}_{\operatorname{Ker}(\bm{X})}\bm{\theta}_{0}^{rl}\|_{2}^{2},$
+    where $\lambda_{1}\geq\cdots\geq\lambda_{r}>0$ are the positive eigenvalues of $(\bm{X}^{\prime}\bm{X})/n$.
+  3. (v)
+    The mean predictive risk is
+    $\operatorname{MPR}(\hat{\bm{\theta}}_{n}^{rl},\bm{\theta}_{0}^{rl})=\frac{r\sigma^{2}}{n}.$
+  4. (vi)
+    If $\operatorname{Range}(\bm{X}^{\prime})=\operatorname{Range}(\bm{\Sigma})$, then $r=r_{0}$ and $\operatorname{Ker}(\bm{X})=\operatorname{Ker}(\bm{\Sigma})$, hence
+    $\operatorname{MSE}(\hat{\bm{\theta}}_{n}^{rl},\bm{\theta}_{0}^{rl})=\frac{\sigma^{2}}{n}\sum_{j=1}^{r_{0}}\frac{1}{\lambda_{j}}\leq\frac{\sigma^{2}r_{0}}{\lambda_{r_{0}}n},\qquad\operatorname{MPR}(\hat{\bm{\theta}}_{n}^{rl},\bm{\theta}_{0}^{rl})=\frac{r_{0}\sigma^{2}}{n}.$
+Proof.
+  1. (i)
+    The model $y=\bm{x}^{\prime}\bm{\theta}_{0}+\varepsilon_{0}$ with $\mathbb{E}[\varepsilon_{0}\mid\bm{x}]=0$ can equivalently be written as
+    $y=\bm{x}^{\prime}\bm{\theta}_{0}^{rl}+\varepsilon(\bm{\theta}_{0}^{rl}),\qquad\mathbb{E}[\varepsilon(\bm{\theta}_{0}^{rl})\mid\bm{x}]=0;$
+    see Proposition 1.1.8. Since $\hat{\bm{\theta}}^{rl}=\bm{X}^{+}\bm{y}$, we have
+    $\mathbb{E}[\hat{\bm{\theta}}^{rl}]=\bm{X}^{+}\mathbb{E}[\bm{y}]=\bm{X}^{+}\bm{X}\bm{\theta}_{0}^{rl}=\bm{P}_{\operatorname{Range}(\bm{X}^{\prime})}\bm{\theta}_{0}^{rl},$
+    which proves the first claim.
+Unbiasedness under $\operatorname{Range}(\bm{X}^{\prime})=\operatorname{Range}(\bm{\Sigma})$ then follows because $\bm{\theta}_{0}^{rl}\in\operatorname{Range}(\bm{\Sigma})=\operatorname{Range}(\bm{X}^{\prime})$.
+  2. (ii)
+    The closed-form expression of $\hat{\bm{\theta}}_{n}^{rl}$ immediately gives
+    $\operatorname{\mathbb{V}ar}[\hat{\bm{\theta}}_{n}^{rl}]=\bm{X}^{+}\operatorname{\mathbb{V}ar}[\bm{\varepsilon}(\bm{\theta}_{0}^{rl})](\bm{X}^{+})^{\prime}.$
+    Since $\bm{\varepsilon}(\bm{\theta}_{0}^{rl})=\bm{X}(\bm{\theta}_{0}-\bm{\theta}_{0}^{rl})+\bm{\varepsilon}_{0}$ and the first term is deterministic given $\bm{X}$, it follows that $\operatorname{\mathbb{V}ar}[\bm{\varepsilon}(\bm{\theta}_{0}^{rl})]=\operatorname{\mathbb{V}ar}[\bm{\varepsilon}_{0}]$.
+  3. (iii)
+    It follows from (ii) since $\bm{X}^{+}(\bm{X}^{+})^{\prime}=(\bm{X}^{\prime}\bm{X})^{+}$.
+  4. (iv)
+    Using $\operatorname{Rank}(\bm{X})=r$, we obtain
+    $\operatorname{tr}(\operatorname{\mathbb{V}ar}[\hat{\bm{\theta}}_{n}^{rl}])=\frac{\sigma^{2}}{n}\operatorname{tr}\bigl((\bm{X}^{\prime}\bm{X}/n)^{+}\bigr)=\frac{\sigma^{2}}{n}\sum_{j=1}^{r}\frac{1}{\lambda_{j}}.$
+    Moreover,
+    $\operatorname{Bias}(\hat{\bm{\theta}}_{n}^{rl},\bm{\theta}_{0}^{rl})=(\bm{P}_{\operatorname{Range}(\bm{X}^{\prime})}-\bm{I})\bm{\theta}_{0}^{rl}=-\bm{P}_{\operatorname{Ker}(\bm{X})}\bm{\theta}_{0}^{rl}.$
+    The result then follows from the bias–variance decomposition in Proposition 1.0.3.
+  5. (v)
+    By Proposition 1.0.1 and $\mathbb{E}[\hat{\bm{\theta}}_{n}^{rl}]=\bm{P}_{\operatorname{Range}(\bm{X}^{\prime})}\bm{\theta}_{0}^{rl}$, we have $\operatorname{lm}(\bm{X},\bm{\theta}_{0}^{rl})=\bm{X}\mathbb{E}[\hat{\bm{\theta}}_{n}^{rl}]$.
+Therefore:
+    $\mathbb{E}[\lVert\operatorname{lm}(\bm{X},\hat{\bm{\theta}}_{n}^{rl})$ $-\operatorname{lm}(\bm{X},\bm{\theta}_{0}^{rl})\rVert_{2}^{2}/n]$
+    $=$ $\mathbb{E}[\lVert\bm{X}(\hat{\bm{\theta}}_{n}^{rl}-\mathbb{E}[\hat{\bm{\theta}}_{n}^{rl}])\rVert_{2}^{2}/n]$
+    $=$ $\mathbb{E}[\operatorname{Trace}\{(\hat{\bm{\theta}}_{n}^{rl}-\mathbb{E}[\hat{\bm{\theta}}_{n}^{rl}])^{\prime}\bm{X}^{\prime}\bm{X}/n(\hat{\bm{\theta}}_{n}^{rl}-\mathbb{E}[\hat{\bm{\theta}}_{n}^{rl}])\}]$
+    $=$ $\mathbb{E}[\operatorname{Trace}\{(\hat{\bm{\theta}}_{n}^{rl}-\mathbb{E}[\hat{\bm{\theta}}_{n}^{rl}])(\hat{\bm{\theta}}_{n}^{rl}-\mathbb{E}[\hat{\bm{\theta}}_{n}^{rl}])^{\prime}\bm{X}^{\prime}\bm{X}/n\}]$
+    $=$ $\operatorname{Trace}(\operatorname{\mathbb{V}ar}[\hat{\bm{\theta}}_{n}^{rl}]\bm{X}^{\prime}\bm{X}/n)$
+    $=$ $\sigma^{2}/n\operatorname{Trace}[(\bm{X}^{\prime}\bm{X})^{+}\bm{X}^{\prime}\bm{X}]=\sigma^{2}/n\operatorname{Trace}(\bm{X}^{+}\bm{X}),$
+    where the last equality follows from the identity $(\bm{X}^{\prime}\bm{X})^{+}\bm{X}^{\prime}=\bm{X}^{+}$.
+    With the spectral decomposition $\bm{X}=\bm{U}\bm{S}\bm{V}^{\prime}$ (Definition 18),
+    $\sigma^{2}/n\operatorname{Trace}(\bm{X}^{+}\bm{X})=$ $\sigma^{2}/n\operatorname{Trace}(\bm{V}(\bm{S}^{\prime})^{+}\bm{S}\bm{V}^{\prime})$
+    $=$ $\sigma^{2}/n\operatorname{Trace}\begin{bmatrix}\bm{I}_{r}&\bm{0}_{r\times(p-r)}\\ \bm{0}_{(p-r)\times r}&\bm{0}_{(p-r)\times(p-r)}\end{bmatrix}$
+    $=$ $\sigma^{2}r/n.$
+  6. (vi)
+    If $\operatorname{Range}(\bm{X}^{\prime})=\operatorname{Range}(\bm{\Sigma})$, then $\operatorname{Rank}(\bm{X})=r_{0}$ and $\operatorname{Ker}(\bm{X})=\operatorname{Ker}(\bm{\Sigma})$.
+Therefore $\operatorname{Bias}(\hat{\bm{\theta}}_{n}^{rl},\bm{\theta}_{0}^{rl})=\bm{0}$ since $\bm{\theta}_{0}^{rl}\in\operatorname{Range}(\bm{\Sigma})$, and thus
+    $\operatorname{MSE}(\hat{\bm{\theta}}_{n}^{rl},\bm{\theta}_{0}^{rl})=\frac{\sigma^{2}}{n}\sum_{j=1}^{r_{0}}\frac{1}{\lambda_{j}}.$
+We now consider the ridge estimator, which introduces an explicit regularization parameter $\lambda>0$ to stabilize the inversion of $\bm{X}^{\prime}\bm{X}$.
+Unlike the ridgeless case, ridge estimation trades bias for reduced variance, yielding a well-defined solution for any $\lambda>0$ and any $(n,p)$ configuration.
+Proposition 1.1.10 (Finite-sample properties of ridge (fixed design)).
+Assume that the linear model $y=\bm{x}^{\prime}\bm{\theta}_{0}+\varepsilon_{0}$ holds for some $\bm{\theta}_{0}\in\mathbb{R}^{p}$ with $\mathbb{E}[\varepsilon_{0}\mid\bm{x}]=0$, and the finite second moments $\mathbb{E}[\|\bm{x}\|_{2}^{2}]<\infty$ and $\mathbb{E}[y^{2}]<\infty$ are met.
+Let $\bm{\Sigma}:=\mathbb{E}[\bm{x}\bm{x}^{\prime}]$ and $r_{0}:=\operatorname{Rank}(\bm{\Sigma})$.
+Fix a realization $\bm{X}\in\mathbb{R}^{n\times p}$ with rows $\bm{x}_{1}^{\prime},\ldots,\bm{x}_{n}^{\prime}$, where each $\bm{x}_{i}$ has the same distribution as $\bm{x}$, and define $\bm{Q}(\lambda):=(\bm{X}^{\prime}\bm{X}+\lambda\bm{I})^{-1}\bm{X}^{\prime}\bm{X}$ for some $\lambda>0$.
+Then (throughout, expectations and variances are understood conditionally on the fixed design $\bm{X}$):
+  1. (i)
+    The ridge estimator is biased:
+    $\mathbb{E}[\hat{\bm{\theta}}_{n}^{r}(\lambda)]=\bm{Q}(\lambda)\bm{\theta}_{0}=\bm{Q}(\lambda)\bm{\theta}_{0}^{rl}.$
+  2. (ii)
+    The variance of the ridge estimator is
+    $\operatorname{\mathbb{V}ar}[\hat{\bm{\theta}}_{n}^{r}(\lambda)]=(\bm{X}^{\prime}\bm{X}+\lambda\bm{I})^{-1}\bm{X}^{\prime}\operatorname{\mathbb{V}ar}[\bm{\varepsilon}(\bm{\theta}_{0}^{rl})]\bm{X}(\bm{X}^{\prime}\bm{X}+\lambda\bm{I})^{-1}.$
+    and $\operatorname{\mathbb{V}ar}[\bm{\varepsilon}(\bm{\theta}_{0}^{rl})]=\operatorname{\mathbb{V}ar}[\bm{\varepsilon}_{0}]$.
+Further assume $\operatorname{\mathbb{V}ar}[\bm{\varepsilon}(\bm{\theta}_{0}^{rl})]=\sigma^{2}\bm{I}$ with $\sigma>0$, and let $r:=\operatorname{Rank}(\bm{X})\leq\min\{n,p\}$. Then:
+  1. (iii)
+    $\operatorname{\mathbb{V}ar}[\hat{\bm{\theta}}_{n}^{r}(\lambda)]=\sigma^{2}(\bm{X}^{\prime}\bm{X}+\lambda\bm{I})^{-1}\bm{X}^{\prime}\bm{X}(\bm{X}^{\prime}\bm{X}+\lambda\bm{I})^{-1}$.
+Moreover, $\operatorname{\mathbb{V}ar}[\hat{\bm{\theta}}_{n}^{rl}]-\operatorname{\mathbb{V}ar}[\hat{\bm{\theta}}_{n}^{r}(\lambda)]$ is positive semidefinite, and it is positive definite on $\operatorname{Range}(\bm{X}^{\prime})$.
+For the remaining statements, assume $\operatorname{Range}(\bm{X}^{\prime})=\operatorname{Range}(\bm{\Sigma})$.
+  1. (iv)
+    The mean squared error is
+    $\operatorname{MSE}(\hat{\bm{\theta}}_{n}^{r}(\lambda),\bm{\theta}_{0}^{rl})=\frac{\sigma^{2}}{n}\sum_{j=1}^{r}\frac{\lambda_{j}}{(\lambda_{j}+\lambda/n)^{2}}+\|[\bm{I}-\bm{Q}(\lambda)]\bm{\theta}_{0}^{rl}\|_{2}^{2},$
+    where $\lambda_{1}\geq\cdots\geq\lambda_{r}>0$ are the positive eigenvalues of $(\bm{X}^{\prime}\bm{X})/n$.
+  2. (v)
+    The mean predictive risk is
+    $\operatorname{MPR}(\hat{\bm{\theta}}_{n}^{r}(\lambda),\bm{\theta}_{0}^{rl})=\frac{\sigma^{2}}{n}\sum_{j=1}^{r}\frac{\lambda_{j}^{2}}{(\lambda_{j}+\lambda/n)^{2}}+\frac{1}{n}\lVert\bm{X}[\bm{Q}(\lambda)-\bm{I}]\bm{\theta}_{0}^{rl}\rVert_{2}^{2}.$
+  3. (vi)
+    Following limits apply:
+    $\lim_{\lambda\to 0}\operatorname{MSE}(\hat{\bm{\theta}}_{n}^{r}(\lambda),\bm{\theta}_{0}^{rl})=\frac{\sigma^{2}}{n}\sum_{j=1}^{r_{0}}\frac{1}{\lambda_{j}}\leq\frac{\sigma^{2}r_{0}}{\lambda_{r_{0}}n},\qquad\lim_{\lambda\to 0}\operatorname{MPR}(\hat{\bm{\theta}}_{n}^{r}(\lambda),\bm{\theta}_{0}^{rl})=\frac{r_{0}\sigma^{2}}{n}.$
+Proof.
+  1. (i)
+    The model $y=\bm{x}^{\prime}\bm{\theta}_{0}+\varepsilon_{0}$ with $\mathbb{E}[\varepsilon_{0}\mid\bm{x}]=0$ can equivalently be written as
+    $y=\bm{x}^{\prime}\bm{\theta}_{0}^{rl}+\varepsilon(\bm{\theta}_{0}^{rl}),\qquad\mathbb{E}[\varepsilon(\bm{\theta}_{0}^{rl})\mid\bm{x}]=0;$
+    see Proposition 1.1.8.
+Since $\hat{\bm{\theta}}_{n}^{r}(\lambda)=(\bm{X}^{\prime}\bm{X}+\lambda\bm{I})^{-1}\bm{X}^{\prime}\bm{y}$ and $\mathbb{E}[\bm{y}]=\bm{X}\bm{\theta}_{0}=\bm{X}\bm{\theta}_{0}^{rl}$, we have
+    $\mathbb{E}[\hat{\bm{\theta}}_{n}^{r}(\lambda)]=(\bm{X}^{\prime}\bm{X}+\lambda\bm{I})^{-1}\bm{X}^{\prime}\bm{X}\bm{\theta}_{0}=\bm{Q}(\lambda)\bm{\theta}_{0},$
+    or
+    $\mathbb{E}[\hat{\bm{\theta}}_{n}^{r}(\lambda)]=(\bm{X}^{\prime}\bm{X}+\lambda\bm{I})^{-1}\bm{X}^{\prime}\bm{X}\bm{\theta}_{0}^{rl}=\bm{Q}(\lambda)\bm{\theta}_{0}^{rl},$
+  2. (ii)
+    From the closed form of $\hat{\bm{\theta}}_{n}^{r}(\lambda)$,
+    $\operatorname{\mathbb{V}ar}[\hat{\bm{\theta}}_{n}^{r}(\lambda)]=(\bm{X}^{\prime}\bm{X}+\lambda\bm{I})^{-1}\bm{X}^{\prime}\operatorname{\mathbb{V}ar}[\bm{\varepsilon}_{0}]\bm{X}(\bm{X}^{\prime}\bm{X}+\lambda\bm{I})^{-1}.$
+  3. (iii)
+    Under $\operatorname{\mathbb{V}ar}[\bm{\varepsilon}_{0}]=\sigma^{2}\bm{I}$,
+    $\operatorname{\mathbb{V}ar}[\hat{\bm{\theta}}_{n}^{r}(\lambda)]=\sigma^{2}(\bm{X}^{\prime}\bm{X}+\lambda\bm{I})^{-1}\bm{X}^{\prime}\bm{X}(\bm{X}^{\prime}\bm{X}+\lambda\bm{I})^{-1}.$
+    Let $\bm{X}=\bm{U}\bm{S}\bm{V}^{\prime}$ with singular values $s_{1}\geq\cdots\geq s_{r}>0$ and write $\lambda_{j}=s_{j}^{2}/n$ (the positive eigenvalues of $(\bm{X}^{\prime}\bm{X})/n$).
+Since $\operatorname{Rank}(\bm{X})=r$, we have $\bm{X}^{\prime}\bm{X}/n=\sum_{j=1}^{r}\lambda_{j}\bm{v}_{j}\bm{v}_{j}^{\prime}$, where $\lambda_{j}=s_{j}^{2}/n$ for $j=1,\ldots,r$.
+It follows $\operatorname{\mathbb{V}ar}[\hat{\bm{\theta}}_{n}^{rl}]=\frac{\sigma^{2}}{n}\sum_{j=1}^{r}\frac{1}{\lambda_{j}}\bm{v}_{j}\bm{v}_{j}^{\prime}$.
+Instead,
+    $\operatorname{\mathbb{V}ar}[\hat{\bm{\theta}}_{n}^{r}(\lambda)]=$ $\sigma^{2}\bm{V}(\bm{S}^{\prime}\bm{S}+\lambda\bm{I})^{-1}\bm{S}^{\prime}\bm{S}(\bm{S}^{\prime}\bm{S}+\lambda\bm{I})^{-1}\bm{V}^{\prime}$
+    $=$ $\frac{\sigma^{2}}{n}\sum_{j=1}^{r}\frac{\lambda_{j}}{(\lambda_{j}+\lambda/n)^{2}}\bm{v}_{j}\bm{v}_{j}^{\prime}.$
+    For each $j\in\{1,\ldots,r\}$:
+    $\frac{1}{\lambda_{j}}-\frac{\lambda_{j}}{(\lambda_{j}+\lambda/n)^{2}}=\frac{(\lambda_{j}+\lambda/n)^{2}-\lambda_{j}^{2}}{\lambda_{j}(\lambda_{j}+\lambda/n)^{2}}=\frac{2\lambda_{j}(\lambda/n)+(\lambda/n)^{2}}{\lambda_{j}(\lambda_{j}+\lambda/n)^{2}}>0.$
+    Therefore the difference is positive semidefinite, and since the coefficients are strictly positive on the span of $\{\bm{v}_{1},\ldots,\bm{v}_{r}\}=\operatorname{Range}(\bm{X}^{\prime})$, it is positive definite on $\operatorname{Range}(\bm{X}^{\prime})$.
+  4. (iv)
+    Using the linearity of the $\operatorname{Trace}$ and the fact that $\bm{V}$ is orthogonal:
+    $\operatorname{Trace}(\operatorname{\mathbb{V}ar}[\hat{\bm{\theta}}_{n}^{r}(\lambda)])=$ $\operatorname{Trace}\left(\frac{\sigma^{2}}{n}\sum_{j=1}^{r}\frac{\lambda_{j}}{(\lambda_{j}+\lambda/n)^{2}}\bm{v}_{j}\bm{v}_{j}^{\prime}\right)$
+    $=$ $\frac{\sigma^{2}}{n}\sum_{j=1}^{r}\frac{\lambda_{j}}{(\lambda_{j}+\lambda/n)^{2}}.$
+    Moreover, if $\operatorname{Range}(\bm{X}^{\prime})=\operatorname{Range}(\bm{\Sigma})$, then
+    $\operatorname{Bias}(\hat{\bm{\theta}}_{n}^{r}(\lambda),\bm{\theta}_{0}^{rl})=[\bm{Q}(\lambda)-\bm{I}]\bm{\theta}_{0}^{rl}.$
+    The result then follows using Proposition 1.0.3.
+  5. (v)
+    Under $\operatorname{Range}(\bm{X}^{\prime})=\operatorname{Range}(\bm{\Sigma})$, we have $\mathbb{E}[\hat{\bm{\theta}}_{n}^{rl}]=\bm{\theta}_{0}^{rl}$ and $\operatorname{\mathbb{V}ar}[\hat{\bm{\theta}}_{n}^{rl}]=\sigma^{2}(\bm{X}^{\prime}\bm{X})^{+}$.
+By definition,
+    $\operatorname{MPR}(\hat{\bm{\theta}}_{n}^{r}(\lambda),\bm{\theta}_{0}^{rl})=\frac{1}{n}\mathbb{E}\lVert\bm{X}(\hat{\bm{\theta}}_{n}^{r}(\lambda)-\bm{\theta}_{0}^{rl})\rVert_{2}^{2}.$
+    Write
+    $\hat{\bm{\theta}}_{n}^{r}(\lambda)-\bm{\theta}_{0}^{rl}=\bm{Q}(\lambda)\hat{\bm{\theta}}_{n}^{rl}-\bm{\theta}_{0}^{rl}=[\bm{Q}(\lambda)-\bm{I}]\bm{\theta}_{0}^{rl}+\bm{Q}(\lambda)\big(\hat{\bm{\theta}}_{n}^{rl}-\bm{\theta}_{0}^{rl}\big).$
+    Therefore,
+    $\frac{1}{n}\mathbb{E}\lVert\bm{X}(\hat{\bm{\theta}}_{n}^{r}(\lambda)-\bm{\theta}_{0}^{rl})\rVert_{2}^{2}$ $=\frac{1}{n}\mathbb{E}\lVert\bm{X}[\bm{Q}(\lambda)-\bm{I}]\bm{\theta}_{0}^{rl}+\bm{X}\bm{Q}(\lambda)(\hat{\bm{\theta}}_{n}^{rl}-\bm{\theta}_{0}^{rl})\rVert_{2}^{2}$
+    $=\frac{1}{n}\lVert\bm{X}[\bm{Q}(\lambda)-\bm{I}]\bm{\theta}_{0}^{rl}\rVert_{2}^{2}+\frac{1}{n}\mathbb{E}\lVert\bm{X}\bm{Q}(\lambda)(\hat{\bm{\theta}}_{n}^{rl}-\bm{\theta}_{0}^{rl})\rVert_{2}^{2}$
+    $\quad+\frac{2}{n}\mathbb{E}\Big\{[\bm{X}[\bm{Q}(\lambda)-\bm{I}]\bm{\theta}_{0}^{rl}]^{\prime}[\bm{X}\bm{Q}(\lambda)(\hat{\bm{\theta}}_{n}^{rl}-\bm{\theta}_{0}^{rl})]\Big\}.$
+    Since $\mathbb{E}[\hat{\bm{\theta}}_{n}^{rl}-\bm{\theta}_{0}^{rl}]=\bm{0}$, the cross term vanishes. Hence
+    $\operatorname{MPR}(\hat{\bm{\theta}}_{n}^{r}(\lambda),\bm{\theta}_{0}^{rl})=\frac{1}{n}\lVert\bm{X}[\bm{Q}(\lambda)-\bm{I}]\bm{\theta}_{0}^{rl}\rVert_{2}^{2}+\frac{1}{n}\mathbb{E}\lVert\bm{X}\bm{Q}(\lambda)(\hat{\bm{\theta}}_{n}^{rl}-\bm{\theta}_{0}^{rl})\rVert_{2}^{2}.$
+    For the second term, use $\mathbb{E}\lVert\bm{A}\bm{z}\rVert_{2}^{2}=\operatorname{Trace}(\bm{A}\operatorname{\mathbb{V}ar}[\bm{z}]\bm{A}^{\prime})$ with $\bm{z}=\hat{\bm{\theta}}_{n}^{rl}-\bm{\theta}_{0}^{rl}$, $\bm{A}=\bm{X}\bm{Q}(\lambda)$:
+    $\frac{1}{n}\mathbb{E}\lVert\bm{X}\bm{Q}(\lambda)(\hat{\bm{\theta}}_{n}^{rl}-\bm{\theta}_{0}^{rl})\rVert_{2}^{2}=\frac{1}{n}\operatorname{Trace}\big(\bm{X}\bm{Q}(\lambda)\operatorname{\mathbb{V}ar}[\hat{\bm{\theta}}_{n}^{rl}]\bm{Q}(\lambda)^{\prime}\bm{X}^{\prime}\big).$
+    Since $\operatorname{\mathbb{V}ar}[\hat{\bm{\theta}}_{n}^{rl}]=\sigma^{2}(\bm{X}^{\prime}\bm{X})^{+}$, this equals
+    $\frac{\sigma^{2}}{n}\operatorname{Trace}\big(\bm{X}^{\prime}\bm{X}\ \bm{Q}(\lambda)(\bm{X}^{\prime}\bm{X})^{+}\bm{Q}(\lambda)^{\prime}\big).$
+    Use $\bm{Q}(\lambda)=(\bm{X}^{\prime}\bm{X}+\lambda\bm{I})^{-1}\bm{X}^{\prime}\bm{X}$ and the SVD $\bm{X}=\bm{U}\bm{S}\bm{V}^{\prime}$ with $\lambda_{j}=s_{j}^{2}/n>0$ for $j=1,\ldots,r$ to diagonalize:
+    $\frac{\sigma^{2}}{n}$ $\operatorname{Trace}\big(\bm{X}^{\prime}\bm{X}\ \bm{Q}(\lambda)(\bm{X}^{\prime}\bm{X})^{+}\bm{Q}(\lambda)^{\prime}\big)$
+    $=\frac{\sigma^{2}}{n}\operatorname{Trace}\Big\{\bm{V}\Big(\frac{\bm{S}^{\prime}\bm{S}}{n}\Big)\Big[\Big(\frac{\bm{S}^{\prime}\bm{S}}{n}+\frac{\lambda}{n}\bm{I}\Big)^{-1}\frac{\bm{S}^{\prime}\bm{S}}{n}\Big]\Big(\frac{\bm{S}^{\prime}\bm{S}}{n}\Big)^{+}\Big(\frac{\bm{S}^{\prime}\bm{S}}{n}\Big)\Big(\frac{\bm{S}^{\prime}\bm{S}}{n}+\frac{\lambda}{n}\bm{I}\Big)^{-1}\bm{V}^{\prime}\Big\}$
+    $=\frac{\sigma^{2}}{n}\sum_{j=1}^{r}\frac{\lambda_{j}^{2}}{(\lambda_{j}+\lambda/n)^{2}}.$
+    Combining the variance and bias contributions yields
+    $\operatorname{MPR}(\hat{\bm{\theta}}_{n}^{r}(\lambda),\bm{\theta}_{0}^{rl})=\frac{\sigma^{2}}{n}\sum_{j=1}^{r}\frac{\lambda_{j}^{2}}{(\lambda_{j}+\lambda/n)^{2}}+\frac{1}{n}\lVert\bm{X}[\bm{Q}(\lambda)-\bm{I}]\bm{\theta}_{0}^{rl}\rVert_{2}^{2},$
+    as claimed.
+  6. (vi)
+    If $\operatorname{Range}(\bm{X}^{\prime})=\operatorname{Range}(\bm{\Sigma})$, then $r=r_{0}$ and $\operatorname{Ker}(\bm{X})=\operatorname{Ker}(\bm{\Sigma})$.
+Moreover, $\bm{Q}(\lambda)=(\bm{X}^{\prime}\bm{X}+\lambda\bm{I})^{-1}\bm{X}^{\prime}\bm{X}\to\bm{\operatorname{P}}_{\operatorname{Range}(\bm{X}^{\prime})}$ as $\lambda\to 0$.
+Since $\bm{\theta}_{0}^{rl}\in\operatorname{Range}(\bm{\Sigma})=\operatorname{Range}(\bm{X}^{\prime})$, the bias term $\|[\bm{I}-\bm{Q}(\lambda)]\bm{\theta}_{0}^{rl}\|_{2}^{2}\to 0$.
+Using the spectral representation $(\bm{X}^{\prime}\bm{X})/n=\sum_{j=1}^{r_{0}}\lambda_{j}\bm{v}_{j}\bm{v}_{j}^{\prime}$ with $\lambda_{j}>0$, we also have
+    $\frac{\sigma^{2}}{n}\sum_{j=1}^{r_{0}}\frac{\lambda_{j}}{(\lambda_{j}+\lambda/n)^{2}}\;\xrightarrow[\lambda\to 0]{}\;\frac{\sigma^{2}}{n}\sum_{j=1}^{r_{0}}\frac{1}{\lambda_{j}}.$
+    Hence
+    $\lim_{\lambda\to 0}\operatorname{MSE}(\hat{\bm{\theta}}_{n}^{r}(\lambda),\bm{\theta}_{0}^{rl})=\frac{\sigma^{2}}{n}\sum_{j=1}^{r_{0}}\frac{1}{\lambda_{j}}\leq\frac{\sigma^{2}r_{0}}{\lambda_{r_{0}}n},\qquad\lim_{\lambda\to 0}\operatorname{MPR}(\hat{\bm{\theta}}_{n}^{r}(\lambda),\bm{\theta}_{0}^{rl})=\frac{r_{0}\sigma^{2}}{n}.$
+The next proposition shows that there are penalty parameter values for which the MSE of ridge is lower than the MSE of ridgeless.
+Proposition 1.1.11.
+Assume that the linear model $y=\bm{x}^{\prime}\bm{\theta}_{0}+\varepsilon_{0}$ holds with $\mathbb{E}[\bm{x}\varepsilon_{0}]=\bm{0}$ and $\operatorname{\mathbb{V}ar}[\bm{\varepsilon}_{0}]=\sigma^{2}\bm{I}$ for $\sigma>0$.
+Then, for a fixed design matrix $\bm{X}$, there exists $\lambda^{*}>0$ such that
+$\operatorname{MSE}(\hat{\bm{\theta}}_{n}^{r}(\lambda^{*}),\bm{\theta}_{0}^{rl})<\operatorname{MSE}(\hat{\bm{\theta}}_{n}^{rl},\bm{\theta}_{0}^{rl}).$
+Proof.
+See Farebrother (1976). ∎
+1.1.6 Finite sample properties of lasso
+In this section, we study the finite sample properties of the lasso estimator under a fixed design matrix $\bm{X}$.
+Since the lasso lacks a closed-form expression, its bias, variance, MSE, and MPR cannot be computed in closed-form as for least squares, ridge, or ridgeless estimators.
+For this reason, rather than attempting explicit moment calculations, we derive non-asymptotic concentration inequalities for its estimation risk $\lVert\hat{\bm{\theta}}_{n}^{l}(\lambda)-\bm{\theta}_{0}\rVert_{2}^{2}$ and predictive risk $\lVert\bm{X}(\hat{\bm{\theta}}_{n}^{l}(\lambda)-\bm{\theta}_{0})\rVert_{2}^{2}/n$.
+To proceed, we first show some auxiliary properties satisfied by any lasso solution.
+Given a set $C$, let $|C|$ denote its cardinality, i.e., the number of elements in $C$.
+For an index set $S\subset\{1,\ldots,p\}$ with complement $S^{c}=\{1,\ldots,p\}\setminus S$, we write $\bm{v}_{S}=[v_{i}]_{i\in S}\in\mathbb{R}^{|S|}$ for the subvector of $\bm{v}\in\mathbb{R}^{p}$ containing the entries indexed by $S$.
+For a constant $\alpha\geq 1$, define the set
+$\mathcal{C}_{\alpha}(S):=\{\bm{v}\in\mathbb{R}^{p}\ :\ \left\lVert\bm{v}_{S^{c}}\right\rVert_{1}\leq\alpha\left\lVert\bm{v}_{S}\right\rVert_{1}\},$
+i.e., the set of vectors in $\mathbb{R}^{p}$ whose coordinates outside $S$ have a total $l_{1}$-size at most $\alpha$ times the $l_{1}$-size of the coordinates within $S$.
+Finally, consider the following definition.
+Definition 20 (Support of a vector).
+The support of vector $\bm{\theta}\in\mathbb{R}^{p}$ is defined as
+$\operatorname{Supp}(\bm{\theta}):=\{j\in\{1,\ldots,p\}\ :\ \theta_{j}\neq 0\}.$
+The next lemma shows that, for a suitable choice of the penalty parameter, the lasso estimator satisfies two key inequalities and that its estimation error lies in $\mathcal{C}_{\alpha}(S)$ for some $\alpha\geq 1$ and an appropriate index set $S$.
+Lemma 1.1.5 (Auxiliary properties of lasso).
+Suppose that the linear model (1.1) holds, and let $s_{0}:=|\operatorname{Supp}(\bm{\theta}_{0})|$.
+If $\lambda\geq 2\left\lVert\bm{X}^{\prime}\bm{\varepsilon}_{0}/n\right\rVert_{\infty}>0$, then any lasso solution $\hat{\bm{\theta}}_{n}^{l}(\lambda)$ satisfies:
+  1. (i)
+    The predictive risk bound
+    $\operatorname{PR}(\hat{\bm{\theta}}_{n}^{l}(\lambda),\bm{\theta}_{0})\leq 12\lambda\left\lVert\bm{\theta}_{0}\right\rVert_{1}.$ (1.19)
+  2. (ii)
+    The estimation error $\hat{\bm{\eta}}:=\hat{\bm{\theta}}_{n}^{l}(\lambda)-\bm{\theta}_{0}$ belongs to $\mathcal{C}_{3}(\operatorname{Supp}(\bm{\theta}_{0}))$ and satisfies
+    $\lVert\bm{X}\hat{\bm{\eta}}\rVert_{2}^{2}/n\leq 3\sqrt{s_{0}}\lambda\lVert\hat{\bm{\eta}}\rVert_{2}.$ (1.20)
+Proof.
+Under the linear model (1.1), we have for any $\bm{\theta}\in\mathbb{R}^{p}$:
+$\left\lVert\bm{y}-\bm{X}\bm{\theta}\right\rVert_{2}^{2}=\bm{y}^{\prime}\bm{y}+\bm{\theta}^{\prime}\bm{X}^{\prime}\bm{X}\bm{\theta}-2\bm{\theta}_{0}^{\prime}\bm{X}^{\prime}\bm{X}\bm{\theta}-2\bm{\varepsilon}_{0}^{\prime}\bm{X}\bm{\theta}.$
+Since $\hat{\bm{\theta}}_{n}^{l}(\lambda)$ is a lasso solution,
+$0\leq\frac{1}{2n}\lVert\bm{y}-\bm{X}\hat{\bm{\theta}}_{n}^{l}(\lambda)\rVert_{2}^{2}+\lambda\lVert\hat{\bm{\theta}}_{n}^{l}(\lambda)\rVert_{1}\leq\frac{1}{2n}\left\lVert\bm{y}-\bm{X}\bm{\theta}_{0}\right\rVert_{2}^{2}+\lambda\left\lVert\bm{\theta}_{0}\right\rVert_{1},$
+which holds if and only if
+$\frac{1}{2n}\lVert\bm{X}\hat{\bm{\eta}}\rVert_{2}^{2}\leq\bm{\varepsilon}_{0}^{\prime}\bm{X}\hat{\bm{\eta}}/n+\lambda(\left\lVert\bm{\theta}_{0}\right\rVert_{1}-\lVert\hat{\bm{\theta}}_{n}^{l}(\lambda)\rVert_{1}).$ (1.21)
+  1. (i)
+    By Hölder inequality,
+    $\bm{\varepsilon}_{0}^{\prime}\bm{X}\hat{\bm{\eta}}/n\leq\lvert\bm{\varepsilon}_{0}^{\prime}\bm{X}\hat{\bm{\eta}}/n\rvert\leq\left\lVert\bm{X}^{\prime}\bm{\varepsilon}_{0}/n\right\rVert_{\infty}\lVert\hat{\bm{\eta}}\rVert_{1}.$
+    Thus, using the choice $\lambda\geq 2\left\lVert\bm{X}^{\prime}\bm{\varepsilon}_{0}/n\right\rVert_{\infty}$ in (1.21) yields
+    $0\leq\frac{1}{2n}\lVert\bm{X}\hat{\bm{\eta}}\rVert_{2}^{2}\leq\lambda/2\lVert\hat{\bm{\eta}}\rVert_{1}+\lambda(\left\lVert\bm{\theta}_{0}\right\rVert_{1}-\lVert\hat{\bm{\theta}}_{n}^{l}(\lambda)\rVert_{1}).$ (1.22)
+    Using the triangle inequality
+    $\lVert\hat{\bm{\eta}}\rVert_{1}\leq\lVert\hat{\bm{\theta}}_{n}^{l}(\lambda)\rVert_{1}+\lVert\bm{\theta}_{0}\rVert_{1},$ (1.23)
+    we further obtain
+    $0\leq\lambda/2(\lVert\hat{\bm{\theta}}_{n}^{l}(\lambda)\rVert_{1}+\lVert\bm{\theta}_{0}\rVert_{1})+\lambda(\lVert\bm{\theta}_{0}\rVert_{1}-\lVert\hat{\bm{\theta}}_{n}^{l}(\lambda)\rVert_{1}),$
+    which, using that $\lambda>0$, implies $\lVert\hat{\bm{\theta}}_{n}^{l}(\lambda)\rVert_{1}\leq 3\lVert\bm{\theta}_{0}\rVert_{1}$.
+Substituting this result into the triangle inequality (1.23) yields
+    $\lVert\hat{\bm{\eta}}\rVert_{1}\leq 4\lVert\bm{\theta}_{0}\rVert_{1}.$ (1.24)
+    Moreover, again by the triangle inequality,
+    $\left\lVert\bm{\theta}_{0}\right\rVert_{1}=\lVert(\bm{\theta}_{0}+\hat{\bm{\eta}})-\hat{\bm{\eta}}\rVert_{1}\leq\lVert\bm{\theta}_{0}+\hat{\bm{\eta}}\rVert_{1}+\lVert\hat{\bm{\eta}}\rVert_{1},$
+    which implies:
+    $\lVert\bm{\theta}_{0}+\hat{\bm{\eta}}\rVert_{1}\geq\left\lVert\bm{\theta}_{0}\right\rVert_{1}-\lVert\hat{\bm{\eta}}\rVert_{1}.$ (1.25)
+    Using (1.24) and (1.25) in the basic inequality (1.22), we obtain
+    $\lVert\bm{X}\hat{\bm{\eta}}\rVert_{2}^{2}/n\leq$ $\lambda\lVert\hat{\bm{\eta}}\rVert_{1}+2\lambda(\left\lVert\bm{\theta}_{0}\right\rVert_{1}-\lVert\bm{\theta}_{0}+\hat{\bm{\eta}}\rVert_{1})$
+    $\leq$ $3\lambda\lVert\hat{\bm{\eta}}\rVert_{1}\leq 12\lambda\lVert\bm{\theta}_{0}\rVert_{1}.$
+  2. (ii)
+    Let $S_{0}:=\operatorname{Supp}(\bm{\theta}_{0})$. Using that $\bm{\theta}_{0S_{0}^{c}}=\bm{0}$, we have
+    $\left\lVert\bm{\theta}_{0}\right\rVert_{1}-\lVert\hat{\bm{\theta}}_{n}^{l}(\lambda)\rVert_{1}=\left\lVert\bm{\theta}_{0S_{0}}\right\rVert_{1}-\lVert\bm{\theta}_{0S_{0}}+\hat{\bm{\eta}}_{S_{0}}\rVert_{1}-\lVert\hat{\bm{\eta}}_{S_{0}^{c}}\rVert_{1}.$
+    Substituting this result into the basic inequality (1.22) yields:
+    $0\leq\lVert\bm{X}\hat{\bm{\eta}}\rVert_{2}^{2}/n\leq\lambda\lVert\hat{\bm{\eta}}\rVert_{1}+2\lambda(\left\lVert\bm{\theta}_{0S_{0}}\right\rVert_{1}-\lVert\bm{\theta}_{0S_{0}}+\hat{\bm{\eta}}_{S_{0}}\rVert_{1}-\lVert\hat{\bm{\eta}}_{S_{0}^{c}}\rVert_{1}).$ (1.26)
+    By the triangle inequality,
+    $\left\lVert\bm{\theta}_{0S_{0}}\right\rVert_{1}=\lVert\bm{\theta}_{0S_{0}}+\hat{\bm{\eta}}_{S_{0}}-\hat{\bm{\eta}}_{S_{0}}\rVert_{1}\leq\lVert\bm{\theta}_{0S_{0}}+\hat{\bm{\eta}}_{S_{0}}\rVert_{1}+\lVert\hat{\bm{\eta}}_{S_{0}}\rVert_{1}.$
+    Therefore, using the decomposition $\lVert\hat{\bm{\eta}}\rVert_{1}=\lVert\hat{\bm{\eta}}_{S_{0}}\rVert_{1}+\lVert\hat{\bm{\eta}}_{S_{0}^{c}}\rVert_{1}$, (1.26) reads
+    $0\leq\lVert\bm{X}\hat{\bm{\eta}}\rVert_{2}^{2}/n\leq$ $\lambda\lVert\hat{\bm{\eta}}\rVert_{1}+2\lambda(\lVert\hat{\bm{\eta}}_{S_{0}}\rVert_{1}-\lVert\hat{\bm{\eta}}_{S_{0}^{c}}\rVert_{1})$
+    $=$ $\lambda(3\lVert\hat{\bm{\eta}}_{S_{0}}\rVert_{1}-\lVert\hat{\bm{\eta}}_{S_{0}^{c}}\rVert_{1}),$
+    which implies that $\hat{\bm{\eta}}\in\mathcal{C}_{3}(\operatorname{Supp}(\bm{\theta}_{0}))$.
+Finally, using the relation between the $l_{1}$- and the $l_{2}$-norm ($\left\lVert\bm{v}\right\rVert_{1}\leq\sqrt{s}\left\lVert\bm{v}\right\rVert_{2}$ for every $\bm{v}\in\mathbb{R}^{s}$), and the notation $s_{0}:=|\operatorname{Supp}(\bm{\theta}_{0})|$, we conclude that
+    $\lVert\bm{X}\hat{\bm{\eta}}\rVert_{2}^{2}/n\leq$ $\lambda(3\lVert\hat{\bm{\eta}}_{S_{0}}\rVert_{1}-\lVert\hat{\bm{\eta}}_{S_{0}^{c}}\rVert_{1})\leq 3\lambda\lVert\hat{\bm{\eta}}_{S_{0}}\rVert_{1}\leq 3\sqrt{s_{0}}\lambda\lVert\hat{\bm{\eta}}_{S_{0}}\rVert_{2}$
+    $\leq$ $3\sqrt{s_{0}}\lambda\lVert\hat{\bm{\eta}}\rVert_{2}.$
+We now derive the main finite-sample properties of lasso under a restricted eigenvalue condition on the design matrix.
+This condition is tailored to the fact—established in Lemma 1.1.5—that for any penalty level $\lambda\geq 2\left\lVert\bm{X}^{\prime}\bm{\varepsilon}_{0}/n\right\rVert_{\infty}$, the estimation error $\hat{\bm{\theta}}_{n}^{l}(\lambda)-\bm{\theta}_{0}$ lies in the cone $\mathcal{C}_{3}(\operatorname{Supp}(\bm{\theta}_{0}))$.
+We therefore impose a lower eigenvalue bound only on vectors in this cone.
+Assumption 1 (Restricted eigenvalue condition).
+The design matrix $\bm{X}\in\mathbb{R}^{n\times p}$ is such that for all $\bm{\eta}\in\mathcal{C}_{3}(\operatorname{Supp}(\bm{\theta}_{0}))$ there exists $\kappa>0$ for which:
+$\left\lVert\bm{X}\bm{\eta}\right\rVert_{2}^{2}/n\geq\kappa\left\lVert\bm{\eta}\right\rVert_{2}^{2},$
+where $\bm{\theta}_{0}\in\mathbb{R}^{p}$ is the coefficient of linear model (1.1).
+In the next proposition we derive finite-sample bounds on the squared $l_{2}$ estimation risk and predictive risk.
+These bounds rely on the restricted eigenvalue condition, which controls the behavior of the design matrix on the cone where the estimation error is known to lie.
+Theorem 1.1.6.
+Suppose that the linear model (1.1) holds and that Assumption 1 is satisfied.
+Let $s_{0}:=|\operatorname{Supp}(\bm{\theta}_{0})|\leq p$.
+Then, any lasso solution $\hat{\bm{\theta}}_{n}^{l}(\lambda)$ with $\lambda\geq 2\left\lVert\bm{X}^{\prime}\bm{\varepsilon}_{0}/n\right\rVert_{\infty}>0$ satisfies:
+  1. (i)
+    The estimation risk bound
+    $\lVert\hat{\bm{\theta}}_{n}^{l}(\lambda)-\bm{\theta}_{0}\rVert_{2}^{2}\leq\frac{9}{\kappa^{2}}s_{0}\lambda^{2}.$ (1.27)
+  2. (ii)
+    The predictive risk bound
+    $\operatorname{PR}(\hat{\bm{\theta}}_{n}^{l}(\lambda),\bm{\theta}_{0})\leq\frac{9}{\kappa}s_{0}\lambda^{2}.$ (1.28)
+Proof.
+In Lemma 1.1.5 we obtained Inequality (1.20), namely
+$\lVert\bm{X}\hat{\bm{\eta}}\rVert_{2}^{2}/n\leq 3\sqrt{s_{0}}\lambda\lVert\hat{\bm{\eta}}\rVert_{2},$
+where $\hat{\bm{\eta}}:=\hat{\bm{\theta}}_{n}^{l}(\lambda)-\bm{\theta}_{0}\in\mathcal{C}_{3}(\operatorname{Supp}(\bm{\theta}_{0}))$.
+  1. (i)
+    Applying Assumption 1 to the left-hand side of the inequality above yields
+    $\kappa\lVert\hat{\bm{\eta}}\rVert_{2}^{2}\leq 3\sqrt{s_{0}}\lambda\lVert\hat{\bm{\eta}}\rVert_{2}.$
+    If $\lVert\hat{\bm{\eta}}\rVert_{2}>0$, dividing both sides by $\lVert\hat{\bm{\eta}}\rVert_{2}$ gives
+    $\lVert\hat{\bm{\eta}}\rVert_{2}\leq 3\sqrt{s_{0}}\lambda/\kappa,$
+    and squaring produces the stated bound. If instead $\lVert\hat{\bm{\eta}}\rVert_{2}=0$, the inequality holds trivially.
+  2. (ii)
+    Applying Assumption 1 to the factor $\lVert\hat{\bm{\eta}}\rVert_{2}$ on the right-hand side of (1.20) yields
+    $\lVert\bm{X}\hat{\bm{\eta}}\rVert_{2}^{2}/n\leq 3\sqrt{s_{0}}\lambda\lVert\bm{X}\hat{\bm{\eta}}\rVert_{2}/\sqrt{n\kappa}.$
+    If $\lVert\bm{X}\hat{\bm{\eta}}\rVert_{2}>0$, dividing both sides by $\lVert\bm{X}\hat{\bm{\eta}}\rVert_{2}/\sqrt{n}$ gives
+    $\lVert\bm{X}\hat{\bm{\eta}}\rVert_{2}/\sqrt{n}\leq 3\sqrt{s_{0}}\lambda/\sqrt{\kappa},$
+    and squaring yields the stated predictive risk bound. If $\lVert\bm{X}\hat{\bm{\eta}}\rVert_{2}=0$, the claim again holds trivially.
+The sparsity parameter $s_{0}=|\operatorname{Supp}(\bm{\theta}_{0})|$ plays a central role in the bounds of Theorem 1.1.6.
+We say that $\bm{\theta}_{0}$ is hard sparse if it contains zero entries.
+More formally:
+Definition 21 (Hard sparsity).
+Coefficient $\bm{\theta}_{0}\in\mathbb{R}^{p}$ is hard sparse if $s_{0}:=|\operatorname{Supp}(\bm{\theta}_{0})|<p$.
+In high-dimensional settings, hard sparsity is typically imposed as an identifying restriction for $\bm{\theta}_{0}$. Consider an asymptotic regime in which
+$\lim_{n\to\infty}p/n=K>0,\quad\text{and}\quad\lim_{p\to\infty}s_{0}/p=s_{\infty}\in(0,1].$
+Under this scaling, $s_{0}\to\infty$ as $n\to\infty$.
+Thus, the lasso converges to $\bm{\theta}_{0}$ only if the restricted eigenvalue $\kappa$ and the penalty level $\lambda$ compensate for the divergence of $s_{0}$, i.e., $\lim_{n\to\infty}s_{0}\lambda^{2}/\kappa^{2}=0$.
+Notice however that $2\left\lVert\bm{X}^{\prime}\bm{\varepsilon}_{0}/n\right\rVert_{\infty}$, the lower bound for $\lambda$ in Theorem 1.1.6, is monotone non-decreasing as additional columns are added to $\bm{X}$.
+Moreover, intuition suggests that Assumption 1 with a large $\kappa$ becomes increasingly more restrictive as $p\to\infty$.
+Remark 7.
+The results in Lemma 1.1.5 and Theorem 1.1.6 can be extended under a milder assumption known as weak sparsity, which formalizes the idea that $\bm{\theta}_{0}$ may be well approximated by a hard sparse vector.
+Definition 22 (Weak sparsity).
+Coefficient $\bm{\theta}_{0}\in\mathbb{R}^{p}$ is weak sparse if $\bm{\theta}_{0}\in B_{q}(r)$ where, for some $q\in[0,1]$ and radius $r>0$,
+$B_{q}(r):=\{\bm{\theta}\in\mathbb{R}^{p}\ :\ \left\lVert\bm{\theta}\right\rVert_{q}^{q}\leq r\}.$
+Setting $q=0$ in Definition 22 recovers hard sparsity with $s=r$.
+Figure 1.10 represents the borders of set $B_{q}(r)$ for $r=1$ and $q=0.5$ and $q=1$.
+For $q\in(0,1]$, weak sparsity restricts the decays of the ordered coefficients: [^7] For $q\in(0,1]$, the map $t\mapsto t^{q}$ is concave, so small coefficients are penalized less than large ones.
+As a consequence, the quantity $\sum_{j=1}^{p}|\theta_{0j}|^{q}$ is dominated by the largest entries.
+For $\bm{\theta}_{0}$ to satisfy $\sum_{j}|\theta_{0j}|^{q}\leq r$, the ordered coefficients $|\theta_{0(1)}|\geq|\theta_{0(2)}|\geq\cdots$ must therefore decay sufficiently quickly.
+$\max_{j=1,\ldots,p}|\theta_{0j}|=\theta_{0(1)}\geq\theta_{0(2)}\geq\ldots,\geq\theta_{0(p-1)}\geq\theta_{0(p)}=\min_{j=1,\ldots,p}|\theta_{0j}|.$
+More precisely, if the coefficients satisfy the bound $|\theta_{0j}|\leq Cj^{-\alpha}$ for some $C\geq 0$ and $\alpha>0$, then $\bm{\theta}_{0}\in B_{q}(r)$ for a radius $r$ depending on $C$ and $\alpha$.
+Figure 1.10: Unit $q$-balls $B_{q}(1)=\{\theta\in\mathbb{R}^{2}:\|\theta\|_{q}^{q}\leq 1\}$ for $q=1$ (blue diamond) and $q=0.5$ (red curve).
+Restricted eigenvalue condition
+Inequality (1.20) of Lemma 1.1.5 provides an upper bound on the predictive risk of the lasso estimator in terms of its estimation error.
+Conversely, for an appropriate choice of the penalty parameter, the restricted eigenvalue condition in Assumption 1 yields an upper bound on the estimation error in terms of predictive risk.
+Combining these two ingredients leads to the estimation and predictive risk bounds in Theorem 1.1.6.
+To provide intuition for why the restricted eigenvalue condition is needed, consider the constrained form of the lasso estimator
+$\hat{\bm{\theta}}_{n}^{lc}(R):=\operatorname*{argmin}_{\bm{\theta}\in\mathbb{R}^{p}}\Bigl\{L_{n}(\bm{\theta}):\|\bm{\theta}\|_{1}\leq R\Bigr\},\qquad R:=\|\bm{\theta}_{0}\|_{1},$
+where
+$L_{n}(\bm{\theta}):=\frac{1}{2n}\|\bm{y}-\bm{X}\bm{\theta}\|_{2}^{2}$
+is the least squares loss.
+With the choice $R=\|\bm{\theta}_{0}\|_{1}$, the true parameter is feasible, and we have $L_{n}(\hat{\bm{\theta}}_{n}^{lc}(R))\leq L_{n}(\bm{\theta}_{0})$.
+Under mild conditions, the loss difference $L_{n}(\bm{\theta}_{0})-L_{n}(\hat{\bm{\theta}}_{n}^{lc}(R))$ decreases with $n$.
+A natural question is when this decrease in the loss translates into a decrease in the estimation error $\|\hat{\bm{\eta}}\|_{2}^{2}$, where $\hat{\bm{\eta}}:=\hat{\bm{\theta}}_{n}^{lc}(R)-\bm{\theta}_{0}$.
+Since $L_{n}$ is quadratic with Hessian $\nabla_{\bm{\theta}}^{2}L_{n}=\bm{X}^{\prime}\bm{X}/n$, the estimation error will shrink when the loss has sufficient curvature in the directions of $\hat{\bm{\eta}}$.
+In a low-dimensional setting ($p\leq n$), a simple sufficient condition ensuring sufficient curvature is that all eigenvalues of $\bm{X}^{\prime}\bm{X}/n$ are uniformly bounded below by a positive constant, that is, for all $\bm{\eta}\in\mathbb{R}^{p}$:
+$\|\bm{X}\bm{\eta}\|_{2}^{2}/n\geq\kappa\|\bm{\eta}\|_{2}^{2}.$
+This global eigenvalue condition is stronger than what is required for the restricted eigenvalue assumption, but it guarantees the desired curvature of the loss function in every direction.
+In the high-dimensional regime ($p>n$), $\bm{X}^{\prime}\bm{X}/n$ is singular and this uniform curvature condition cannot hold globally.
+However, Lemma 1.1.5 shows that the lasso estimation error lies in the cone $\mathcal{C}_{3}(\operatorname{Supp}(\bm{\theta}_{0}))$.
+The restricted eigenvalue condition requires curvature only along these directions.
+Note that this condition does not require $\bm{X}_{S_{0}}$, the matrix formed by the columns of $\bm{X}$ with index in $S_{0}$, to have full column rank, nor does it imply $|S_{0}|\leq\operatorname{Rank}(\bm{X})$.
+It simply demands that the design has sufficient curvature within the cone of directions in which the lasso error can lie.
+Under this restricted curvature condition, a small loss difference still implies a small estimation error, even when $p\gg n$.
+Verifying that a given deterministic design satisfies the restricted eigenvalue condition is challenging.
+Developing probabilistic conditions ensuring that random designs satisfy it with high probability is an active area of research.
+Slow rates and fast rates
+Consider the case in which the error term in the linear model (1.1) is sub-Gaussian with mean zero and variance proxy $\sigma^{2}$.
+Under this assumption, one can select a tuning parameter $\lambda$ that depends only on the unknown $\sigma^{2}$ and guarantees that the estimation and prediction risks of the lasso estimator are upper bounded with high probability.
+Theorem 1.1.7.
+Suppose that the linear model (1.1) holds and that $\bm{\varepsilon}_{0}$ is a vector of independent random variables $\varepsilon_{0i}\sim\mathrm{subG}(\sigma^{2})$ with variance proxy $\sigma^{2}>0$.
+Further suppose that the columns of $\bm{X}$ are standardized so that $\max_{j=1,\ldots,p}\left\lVert\bm{X}_{j}\right\rVert_{2}/\sqrt{n}\leq C$ for some constant $C>0$.
+Then, for all $\delta>0$, any lasso solution $\hat{\bm{\theta}}_{n}^{l}(\lambda)$ with regularization parameter
+$\lambda=2C\sigma\left(\sqrt{2\ln(p)/n}+\delta\right)$ (1.29)
+satisfies with probability at least $1-2e^{-n\delta^{2}/2}$:
+$\operatorname{PR}(\hat{\bm{\theta}}_{n}^{l}(\lambda),\bm{\theta}_{0})\leq 24C\left\lVert\bm{\theta}_{0}\right\rVert_{1}\sigma(\sqrt{2\ln(p)/n}+\delta).$ (1.30)
+Further suppose that Assumption 1 holds and let $s_{0}:=|\operatorname{Supp}(\bm{\theta}_{0})|\leq p$. Then, with probability at least $1-2e^{-n\delta^{2}/2}$:
+  1. (i)
+    The estimation risk bound
+    $\lVert\hat{\bm{\theta}}_{n}^{l}(\lambda)-\bm{\theta}_{0}\rVert_{2}^{2}\leq\frac{72C^{2}\sigma^{2}s_{0}}{\kappa^{2}}(2\ln(p)/n+\delta^{2}).$ (1.31)
+  2. (ii)
+    The predictive risk bound
+    $\operatorname{PR}(\hat{\bm{\theta}}_{n}^{l}(\lambda),\bm{\theta}_{0})\leq\frac{72C^{2}\sigma^{2}s_{0}}{\kappa}(2\ln(p)/n+\delta^{2}).$ (1.32)
+Proof.
+For any $t\geq 0$, by the union bound:
+$\mathbb{P}\left[\left\lVert\bm{X}^{\prime}\bm{\varepsilon}_{0}/n\right\rVert_{\infty}\geq t\right]$ $=\mathbb{P}\left[\max_{j=1,\ldots,p}\lvert\bm{X}_{j}^{\prime}\bm{\varepsilon}_{0}/n\rvert\geq t\right]$
+$=\mathbb{P}\left[\bigcup_{j=1}^{p}\left\{\,\lvert\bm{X}_{j}^{\prime}\bm{\varepsilon}_{0}/n\rvert\geq t\,\right\}\right]$
+$\leq\sum_{j=1}^{p}\mathbb{P}\left[\lvert\bm{X}_{j}^{\prime}\bm{\varepsilon}_{0}/n\rvert\geq t\right],$
+where the second equality follows from the fact that the maximum of a set of numbers exceeds $t$ if and only if at least one element in the set exceeds $t$.
+Since $\varepsilon_{01},\ldots,\varepsilon_{on}$ are independent random variables with $\mathrm{subG}(\sigma^{2})$ distribution, from Proposition LABEL:app:prop:sum_indep_sub_gauss we have that for any $t\in\mathbb{R}$:
+$\sum_{j=1}^{p}\mathbb{P}\left[\lvert\bm{X}_{j}^{\prime}\bm{\varepsilon}_{0}/n\rvert\geq t\right]$ $\leq\sum_{j=1}^{p}2\exp\left(-\frac{t^{2}}{2\sigma^{2}\left\lVert\bm{X}_{j}/n\right\rVert_{2}^{2}}\right)$
+$\leq 2p\exp\left(-\frac{t^{2}n}{2\sigma^{2}C^{2}}\right),$
+where the last inequality follows from the standardization of the columns of $\bm{X}$, namely
+$\left\|\bm{X}_{j}/n\right\|_{2}^{2}=\|\bm{X}_{j}\|_{2}^{2}/n^{2}\leq C^{2}n/n^{2}=C^{2}/n.$
+Substituting $t=C\sigma\left(\sqrt{2\ln(p)/n}+\delta\right)$ we get
+$2p\exp\left(-\frac{t^{2}n}{2\sigma^{2}C^{2}}\right)$ $=2\exp(\ln(p))\exp\left(-n\delta^{2}/2-\ln(p)-\delta\sqrt{2n\ln(p)}\right)$
+$=2\exp\left(-n\delta^{2}/2\right)\exp\left(-\delta\sqrt{2n\ln(p)}\right)$
+$\leq 2\exp\left(-n\delta^{2}/2\right),$
+since $\delta\sqrt{2n\ln(p)}<0$. We conclude that, for all $\delta>0$:
+$\mathbb{P}[2\left\lVert\bm{X}^{\prime}\bm{\varepsilon}_{0}/n\right\rVert_{\infty}\leq 2C\sigma(\sqrt{2\ln(p)/n}+\delta)]\geq 1-2e^{-n\delta^{2}/2}.$
+Consequently, if we set $\lambda=2C\sigma(\sqrt{2\ln(p)/n}+\delta)$, we obtain from (1.19) of Lemma 1.1.5 that (1.30) holds with probability at least $1-2e^{-n\delta^{2}/2}$.
+Moreover, under Assumption 1, we obtain from (1.27) and (1.28) of Theorem 1.1.6 that (1.31) and (1.32) hold with probability at least $1-2e^{-n\delta^{2}/2}$, by using the inequality: [^8] This inequality follows from $2ab\leq a^{2}+b^{2}$ for any two real numbers $a$ and $b$.
+$2\ln(p)/n+\delta^{2}+2\sqrt{2\ln(p)/n}\delta\leq 2(2\ln(p)/n+\delta^{2}).$
+As long as $n\geq 2\ln(p)$, the ratio $2\ln(p)/n$ can be much smaller than $\sqrt{2\ln(p)/n}$.
+This distinction motivates the terminology commonly used in the high-dimensional statistics literature.
+The bound (1.30), which scales on the order of
+$\sqrt{2\ln(p)/n},$
+is referred to as the slow rate for the predictive risk of the lasso, whereas the bound (1.32), which scales on the order of
+$2\ln(p)/n,$
+is called the fast rate. These rates are depicted in Figure 1.11.
+The slow rate arises from ensuring that the tuning parameter satisfies $\lambda\geq 2\|\bm{X}^{\prime}\bm{\varepsilon}_{0}/n\|_{\infty}$, a condition that can be enforced with high probability under sub-Gaussian errors and column normalization of the design.
+This argument does not rely on any structural assumptions on $\bm{X}$ beyond these probabilistic and scaling considerations, and therefore the resulting bound holds without the restricted eigenvalue condition.
+The slow rate thus provides a general upper bound on the predictive risk that depends $\|\bm{\theta}_{0}\|_{1}$, $n$ and $p$.
+By contrast, the fast rate requires additional geometric structure on the design, as formalized by the restricted eigenvalue condition.
+This assumption guarantees sufficient curvature of the least squares loss along the cone in which the lasso estimation error lies.
+Under this condition, the lasso achieves sharper bounds of order $s_{0}\ln(p)/n$ up to multiplicative constants.
+In cases where $s_{0}$ is much smaller than $p$, this represents a substantial improvement over the slow rate.
+Figure 1.11: Comparison of slow and fast lasso rates for $n=50$ and $p\in[1,20]$.